Linux Consultants HOWTO

Martin Michlmayr, tbm@cyrius.com, http://www.cyrius.com/tbm v5.25, 11 April 1997 This document contains a listing of companies providing commercial Linux related support. It is maintained by Martin Michlmayr <tbm@cyrius.com>.

Contents

1 About this document 10

2 Copyright Information 10

3 Argentina 11 3.1 CIBER*NET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.2 InFoDaX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

4 Australia 12 4.1 Echo Labs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 4.2 Farrow Norris Pty Ltd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 4.3 Interactive Infotainment Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.4 InterWeft . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.5 IPAX Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 4.6 TereDonn Computer Engineering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

5 Austria 16 5.1 Bergmann Engineering and Consulting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 5.2 Company Systems Consulting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 5.3 g.a.m.s. edv-dienstleistungen gmbh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 5.4 EDV-Beratung Friedrich Ganter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 5.5 Horus GmbH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 5.6 Innovative Data Administration Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 5.7 Alexander List . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 5.8 Thomas Michlmayr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 5.9 netwing GmbH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 5.10 Quant X EDV-Ges.m.b.H. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 6 Belgium 24

6.1 Double Barrel Consultancy and Productions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 6.2 Phidani Software SPRL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

7 Brazil 25 7.1 José Carlos Benfati . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

8 Cambodia 26 8.1 Mike Gärtner . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

9 Canada 27 9.1 Cabala Systems Corp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 9.2 Centurion Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 9.3 Dataden Computer Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 9.4 Datasoft Communications . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 9.5 Dynamicro Consulting Limited . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 9.6 Global Proximity Corporation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 9.7 G.M. Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 9.8 Hard Data Ltd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 9.9 IED Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 9.10 Jones Computer Supplies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 9.11 MostlyLinux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 9.12 Gary Lawrence Murphy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 9.13 Pyx Networking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 9.14 Ken Roberts Computer Consultants Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 9.15 Silver Technologies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 9.16 Troika Management Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

10 Denmark 39 10.1 Lasse Oesterild . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

11 Europe 39 11.1 Jochen Hayek . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 11.2 Tarkus AB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 12 Finland 41

12.1 Netsol Network Solutions Oy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

13 Germany - General 42 13.1 Feichtinger IT Consulting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 13.2 mdt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 13.3 MetaWorks GmbH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 13.4 Ntakpe Management & Consulting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 13.5 S.u.S.E. Gmbh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

14 Germany - Baden-Württemberg 45 14.1 Hans-Joachim Baader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 14.2 ISS Internet Service Stuttgart GmbH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 14.3 MU Softwareentwicklung . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 14.4 Jochen Wiedmann . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 14.5 Michael Will Linux Systeme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

15 Germany - Bayern 49 15.1 B.I.S Babylon Internet Service GmbH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 15.2 EDV-Beratung Gert Döring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 15.3 K&K Kommunikationssysteme GmbH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 15.4 noris network GmbH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 15.5 Julian Howard Stacey, Vector Systems Limited, München . . . . . . . . . . . . . . . . . . . . 52 15.6 WorNet Internetdienste . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

16 Germany - Berlin 53 16.1 netcomp Informationsmanagement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 16.2 Woller & Link GbR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

17 Germany - Brandenburg 55 17.1 Bernd Hentig EDV-Beratung . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

18 Germany - Hamburg 56 18.1 Jens Doll . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 18.2 Dipl.-Ing. Axel Ramge . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 19 Germany - Hessen 57

19.1 Gerd Aschemann . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 19.2 b.i.t. beratungsgesellschaft für informations-technologie mbh . . . . . . . . . . . . . . . . . . 58 19.3 LEMIS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

20 Germany - Mecklenburg-Vorpommern 60 20.1 Bernd Wolf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60

21 Germany - Niedersachsen 60 21.1 Gärtner Datensysteme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 21.2 Jürgen Gerloff Elektronik . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 21.3 Electronic & Computertechnik Margitta Holzbrecher . . . . . . . . . . . . . . . . . . . . . . . 62 21.4 Martin Schulze . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

22 Germany - Nordrhein-Westfalen 63 22.1 Stephan Austermühle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 22.2 dev/consulting GmbH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 22.3 Ing. Büro Gerichhausen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 22.4 netServe oHG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 22.5 PiN - Präsenz im Netz GmbH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 22.6 Silent Running . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 22.7 SUNSOFT Büro und EDV GmbH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 22.8 transdata - Bigge & Langanke GbR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 22.9 Jörg Weber . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

23 Germany - Saarland 70 23.1 B. Neubert EDV-Dienstleistungen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

24 Germany - Schleswig-Holstein 71 24.1 AllCon GmbH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

25 Germany - Thüringen 72 25.1 IKS GmbH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 26 Hungary 73

26.1 Krisztián Czakó . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

27 India 74 27.1 Algologic Research & Solutions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

28 Israel 74 28.1 TclTek Ltd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

29 Italy 75 29.1 P.O.P. Art S.r.l. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 29.2 PROFUSO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 29.3 SIR s.r.l. - PublinetWork . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 29.4 Softmedia Spa . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 29.5 Studio LEADER Pro . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

30 Japan 79 30.1 SoftWare Research Co. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

31 Mexico 80 31.1 Miguel de Icaza . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 31.2 Laboratorio de Sistemas Linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

32 The Netherlands 81 32.1 Berrymount Automation B.V. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 32.2 Bitbybit Information Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 32.3 BitWizard B.V. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 32.4 Cistron Internet Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 32.5 MANTRA Research . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 32.6 Netco . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 32.7 Le Reseau netwerksystemen BV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86

33 Norway 87 33.1 it-analyse as . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 33.2 NorConnect Internet Services AS/NorConnect Group . . . . . . . . . . . . . . . . . . . . . . . 88

33.3 UNISON AS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88

34 Pakistan 89 34.1 Vision Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

35 Poland 90 35.1 Polskie Sieci Teleinformatyczne S.A. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 35.2 Publiczny Dostep do Internetu sp. z o.o. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

36 Singapore 92 36.1 Netstream Pte Ltd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

37 South Africa 92 37.1 Obsidian Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 37.2 Symphony Research (Pty) Ltd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

38 Spain 94 38.1 DISAM-SL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

39 Sweden 95 39.1 Guide Datakonsult . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 39.2 Signum Support AB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96

40 Turkey 96 40.1 Hakan Kalyoncu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96

41 United Kingdom 97 41.1 Cygnet Internet Services Ltd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 41.2 Paul Dunne . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 41.3 Laboratory for Micro Enterprise . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 41.4 Marketgrid Limited . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 41.5 Nene Information Sytems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100

42 USA - General 101 42.1 InfoMagic, Inc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 42.2 NDA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

42.3 PSA AnswerDesk – Pencom Systems Administration . . . . . . . . . . . . . . . . . . . . . . . 103

43 USA - Arizona 103 43.1 M. Cooper . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 43.2 Mechanix Computer Consulting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 43.3 M&S Group, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105

44 USA - California 106 44.1 Evolve Computer Solutions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 44.2 Inter@ctivate Consulting Group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 44.3 Linux-Consulting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 44.4 RCT Design . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 44.5 RG Consulting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 44.6 Bill Rousseau . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 44.7 Softcraft Impresa . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 44.8 Earl A. Stutes, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 44.9 Vennerable Consultants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 44.10Gregg Weber . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 44.11Yggdrasil Computing, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113

45 USA - Connecticut 114 45.1 BRT Technical Services Corporation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 45.2 Creative Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 45.3 Darcom Systems Ltd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 45.4 Kracked Rock Komputing, LLC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116

46 USA - Colorado 117 46.1 Colorado Computer Consultants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 46.2 Daylight Software . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 46.3 Granite Computing Solutions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 46.4 SNLutz Associates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 46.5 Worldwide Solutions, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 47 USA - Florida 120

47.1 LOD Communications . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 47.2 Progressive Computer Concepts, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121

48 USA - Georgia 122 48.1 Fly-By-Day Consulting, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 48.2 Triangle Software Corporation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123

49 USA - Illinois 124 49.1 Rahim Azizarab . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 49.2 onShore, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124

50 USA - Indiana 125 50.1 Citadel Software, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 50.2 LANscape . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126

51 USA - Maine 127 51.1 Midcoast Internet Solutions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127

52 USA - Maryland 127 52.1 I-Link, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 52.2 Matthew Kaylor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 52.3 Bradley M. Kuhn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 52.4 Vertex Development . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130

53 USA - Massachusetts 130 53.1 Working Version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130

54 USA - Michigan 131 54.1 Jim Willette . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131

55 USA - Missouri 132 55.1 Reboot Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 55.2 Damon Richardson . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 56 USA - Nebraska 134

56.1 Probe Technology Inc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 56.2 tummy.com, ltd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134

57 USA - Nevada 135 57.1 Steve Hedlund . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

58 USA - New Jersey 136 58.1 Etc Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 58.2 ZEI Software . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137

59 USA - New York 137 59.1 Crynwr Software . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137

60 USA - Ohio 138 60.1 DESIGN SCIENCE LABS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 60.2 Internet Robotics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 60.3 Practical Network Design . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140

61 USA - Oregon 141 61.1 Kymsoft Consulting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 61.2 NET-Community . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141

62 USA - Pennsylvania 142 62.1 Christopher J. Fearnley . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142

63 USA - Texas 143 63.1 xprt Computer Consuting, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143

64 USA - Virginia 144 64.1 Internet Systems and Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144

65 USA - Washington 144 65.1 Cutting Edge Communications, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144

66 USA - Wisconsin 145 66.1 Couvares Consulting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145

1 About this document

This is the Linux Consultants HOWTO. It is a listing of companies providing commercial Linux related support. It is maintained by Martin Michlmayr <tbm@cyrius.com>.

If you need to know more about the Linux Documentation Project or about Linux HOWTO’s, feel free to contact the supervisor Greg Hankins <gregh@sunsite.unc.edu>.

Greg Hankins will post the listing to several national and international newsgroups on a monthly basis. In addition, the Linux Consultants HOWTO can be found on the World Wide Web at http://www.cyrius.com/tbm/Consultants-HOWTO. New versions of the Linux Consultants HOWTO are always placed at this site first, so please be sure to check if the copy you are reading is still up to date!

Companies providing Linux support are invited to fill out the following form and contact me at tbm@cyrius.com.

     Name:
     Address (If you are located in Germany, please specify your Bundesland):
     Phone:
     Fax:
     EMail:
     URL:
     Contact (Name of contact person):
     Type of support (e.g. phone, EMail, remote network
             administration, inhouse, onsite):
     Special expertise (e.g. installation, programming services, system
             administration, Internet and Intranet connectivity,
             firewall installation, troubleshooting, training):
     Sample prices (please specify currency):

2 Copyright Information

This HOWTO is Copyright \begin{matrix}c \\ ◯ \end{matrix} 1996, 1997 by Martin Michlmayr.

A verbatim copy may be reproduced or distributed in any medium physical or electronic without permission of the author. Translations are similarly permitted without express permission if it includes a notice on who translated it.

Short quotes may be used without prior consent by the author. Derivative work and partial distributions of the Linux Consultants HOWTO must be accompanied with either a verbatim copy of this file or a pointer to the verbatim copy.

Commercial redistribution is allowed and encouraged; however, the author would like to be notified of any such distributions.

In short, we wish to promote dissemination of this information through as many channels as possible. However, we do wish to retain copyright on the HOWTO documents, and would like to be notified of any plans to redistribute the HOWTOs.

We further want that all information provided in the HOWTOs is disseminated. If you have questions, please contact Greg Hankins, the Linux HOWTO coordinator, at gregh@sunsite.unc.edu.

3 Argentina

3.1 CIBER*NET

          Derqui 47
          5501 Godoy Cruz
          Mendoza
          Argentina

Phone:

+54 61 243883 EMail:

afernand@planet.losandes.com.ar Contact:

Alejandro Fernandez Herrero Type of support:

Remote network and onsite. Special expertise:

General Unix and Linux consulting, networking, connectivity with other operating systems, support, porting and web pages development. Caldera support.

Sample prices:

Upon request. Last modified:

March 24, 1997.

3.2 InFoDaX

          Osvaldo Magnasco 771
          1706 Haedo
          Buenos Aires
          Argentina

Phone:

54 1 659 0697 Fax:

+54 1 659 0697 EMail:

ocinfoda@starnet.net.ar Contact:

Oscar Andres Carriles Type of support:

Remote network and onsite. Special expertise:

General Unix and Linux consulting, support, aplications and development. Caldera and FlagShip. Sample prices:

Upon request. Last modified:

March 26, 1997.

4 Australia

4.1 Echo Labs

          29 Weld St
          Nedlands, WA 6009
          Australia

Phone:

+61 9 389 1680, +61 41 985 9603 (mobile) Fax:

+61 9 389 1351 EMail:

echo@opera.iinet.net.au URL:

http://www.iinet.net.au/˜ echo Contact:

Ralph Billes Type of support:

Phone, EMail, remote network administration, inhouse and onsite. Special expertise:

Networking, installations, custom software solutions, system administration, Internet and Intranet connectivity, firewall installation, problem solving.

Sample prices:

AUS$75/hour onsite. Last modified:

March 7, 1997.

4.2 Farrow Norris Pty Ltd

          Epping
          Sydney, NSW 2121
          Australia
          Hurstville
          Sydney, NSW 2220
          Australia

Phone:

+61 41 7243239 (mobile) EMail:

info@fn.com.au URL:

http://www.usyd.edu.au/˜ jfarrow Contact:

Stephen Norris Type of support:

Phone, EMail, remote network administration, inhouse and onsite. Special expertise:

Installations, system administration, training and Unix programming. Sample prices:

AUS$40-$70 (depends on length, complexity). Last modified:

March 16, 1997.

4.3 Interactive Infotainment Systems

          20 Ivanhoe Rd
          Croydon, NSW 2132
          Australia

Phone:

+61 2 9798 7604, +61 15 955 042 (mobile) EMail:

anthony@rumble.waratah.id.au URL:

http://www.infotainment.com.au Contact:

Anthony Rumble Type of support:

Phone, EMail and inhouse. Special expertise:

Training, porting, development, web and CGI, networking. Sample prices:

Upon request. Last modified:

March 15, 1997.

4.4 InterWeft

          Melbourne
          Australia

Phone:

+61 3 9735 3586 Fax:

+61 3 9735 3586 EMail:

info@interweft.com.au URL:

http://www.interweft.com.au Contact:

Robert Hart Type of support:

Phone, EMail, fax, remote network administration and onsite. Special expertise:

Strategic IT business planning. Data/voice networking (design, specification and integration), Internet connectivity planning, design, installation, configuration and security. Linux/Unix installation, configuration, training and support. Author of Linux PPP-HOWTO and mail2news mini-HOWTO.

Sample prices:

Upon request. Quick help via EMail is provided free as long as the work load permits - particularly for Linux PPP.

Last modified:

March 16, 1997.

4.5 IPAX Systems

     GPO Box 94A
     Melbourne, VIC 3001
     Australia

Phone:

+61 3 9887 1984 (Network Operations Centre Phone) Fax:

+61 3 9801 8533 EMail:

info@ipax.com.au URL:

http://www.ipax.com.au Contact:

Adam Neat Type of support:

EMail. Special expertise:

Specialized in system design, systems management and online services. We use Linux for the majority of the Unix machines we install. 24 hour NOC services are also available. Linux is used for this.

Sample prices:

Free for advice. Last modified:

March 16, 1997.

4.6 TereDonn Computer Engineering

          14/357 Logan Road, Stones Corner
          Brisbane, QLD 4120
          Australia

Phone:

+61 7 38474966, +61 414 663053 (24 hours) Fax:

+61 7 38473599 EMail:

terry@tdce.com.au URL:

http://www.tdce.com.au Contact:

Terry Sweetser Type of support:

Remote network administration and onsite. Special expertise:

Internet gateways, firewalls. Sample prices:

AUS$100/hour. Last modified:

March 17, 1997.

5 Austria

5.1 Bergmann Engineering and Consulting

          Dirmhirngasse 27/118
          A-1230 Wien
          Austria

Phone:

+43 1 8695186, +43 2254 7802703, +43 664 3573565 Fax:

+43 1 319956110 EMail:

office@bec.co.at URL:

http://www.bec.co.at Contact:

Josef Bergmann Type of support:

Phone, EMail and inhouse. Special expertise:

Installation, troubleshooting, training, administration, consulting, Internet and Intranet connectivity, Linux file, print, fax, EMail, web, FTP, DNS, cache, modem servers. Custom software development (C and C++) for scientific, industrial or business applications

Sample prices:

Base price ATS 950/hour (approximately US$80). Last modified:

February 28, 1997.

5.2 Company Systems Consulting

     Pohlgasse 41/14
     A-1120 Wien
     Austria

Phone:

+43 222 8170494 AirPage:

+43 688 4220313 EMail:

mike@xphys.tuwien.ac.at URL:

http://sus1.xphys.tuwien.ac.at/mike Contact:

Michael Rumpler Type of support:

Phone, EMail, remote network administration, inhouse and onsite. Special expertise:

Installation, programming services, system administration, Internet and Intranet connectivity, firewall installation, troubleshooting, training, web services, special software solutions.

Sample prices:

ATS 900/hour. Last modified:

March 15, 1997.

5.3 g.a.m.s. edv-dienstleistungen gmbh

     Sch\"{o}nbrunnerstr. 179/II/4
     A-1120 Vienna
     Austria

Phone:

+431 8176230 Fax:

+431 8176230 17 EMail:

linux@gams.at URL:

http://www.gams.at Contact:

Vinzenz Grabner, Michael Lausch, Rainer Hochreiter Type of support:

Phone or EMail for commercial customers. Special expertise:

Linux installation, configuration and training, networking (Internet and Intranet), databases under Linux (developed Solid perl and python interface), security, design of security policies and firewall configurations, integration of heterogenous networks (Appletalk, SMB), development of network management applications (SNMP and TMN), programming of Internet applications (e.g. http://www.gams.at/ubahn), ISP (e.g. leased lines, dial-up, web).

Sample prices:

Upon request. Last modified:

March 23, 1997.

5.4 EDV-Beratung Friedrich Ganter

     Grazerstra{\ss}e 26a
     A-8045 Graz
     Austria

Phone:

+43 316 688112, +43 664 1021621 Fax:

+43 316 688112 4 EMail:

ganter@ganter.at URL:

http://fvkma.tu-graz.ac.at/fritz Contact:

Friedrich Ganter Type of support:

Phone, EMail and inhouse. Special expertise:

Alpha Support only through EDV-Consulting Wolfgang DREYER. Linux installation, turnkey systems for firewall, web and mail server, Intranet server. Choose Intel or Digital Alpha based systems.

Sample prices:

US$40-$80/hour. Last modified:

February 21, 1997.

5.5 Horus GmbH

          Jakob-Haringer-Stra{\ss}e 8
          A-5020 Salzburg
          Austria

Phone:

+43 662 450363 Fax:

+43 662 450364 EMail:

info@horus.com URL:

http://www.horus.com Contact:

Thomas “Dune” Freina Type of support:

EMail and inhouse. Special expertise:

Consulting, installation, training and programming. Sample prices:

Base price is ATS 1250/hour (approximately US$115). Last modified:

March 17, 1997.

5.6 Innovative Data Administration Services

          iDAS-Consulting
          z. Hdn. Thomas Kirchtag
          Gasgasse 13
          A-1150 Wien
          Austria

Phone:

+43 664 2231623 Fax:

+43 1 6898252 EMail:

office@iDAS.co.at URL:

http://www.iDAS.co.at Contact:

Thomas Kirchtag Type of support:

Phone, remote network administration and onsite. Special expertise:

System administration, Internet and Intranet services, web services, client/server database systems, firewalls, secure web server.

Sample prices:

ATS 800/hour. Last modified:

April 7, 1997.

5.7 Alexander List

          Jahngasse 5
          A-8010 Graz
          Austria

Phone:

+43 316 815234 EMail:

alexlist@sbox.tu-graz.ac.at URL:

http://www.sbox.tu-graz.ac.at/home/alexlist Contact:

Alexander List Type of support:

Phone, EMail and inhouse. Special expertise:

Specialized in ISDN solutions.

Sample prices:

Basic rate US$60/hour. To be negotiated, depending on type and duration of job. Last modified:

March 17, 1997.

5.8 Thomas Michlmayr

          Brunnengasse 4
          A-5020 Salzburg
          Austria

Phone:

+43 662 450072 EMail:

mike@backbone.at Contact:

Thomas Michlmayr Type of support:

EMail, remote network administration, inhouse and onsite. Special expertise:

System administration (e.g. DNS, mail, news), software installation and porting, training for system administrators.

Sample prices:

Upon request. Last modified:

March 17, 1997.

5.9 netwing GmbH

          Brixentaler Str. 12
          A-6300 W\"{o}rgl
          Austria

Phone:

+43 5332 70455 Fax:

+43 5332 70455 5 EMail:

info@netwing.at URL:

http://www.netwing.at Contact:

Peter Saringer Type of support:

Phone, EMail and inhouse. Special expertise:

Training for commercial customers. Linux installation and configuration, networking, security policies, firewall consulting and configuration. Internet and Intranet solutions.

Sample prices:

ATS 1100/hour. Last modified:

March 19, 1997.

5.10 Quant X EDV-Ges.m.b.H.

          Hintnausdorf 14/1
          A-9300 St. Veit/Glan
          Austria

Phone:

+43 4212 60040 Fax:

+43 4212 60044 EMail:

support@quant-x.com URL:

http://www.quant-x.at Contact:

Wolfgang Dreyer Type of support:

EMail, phone and onsite.

Special expertise:

Specialized on Linux/Alpha. Sample prices:

30 days getting started US$150. Last modified:

March 15, 1997.

6 Belgium

6.1 Double Barrel Consultancy and Productions

          Sportstraat 28
          9000 Gent
          Belgium

Phone:

+32 9 2227764 Fax:

+32 9 2224976 EMail:

mvergall@double-barrel.be URL:

http://www.double-barrel.be Contact:

Michael C. Vergallen Type of support:

Phone, EMail and inhouse. Special expertise:

Linux installation support, Intranet and Internet setups, web design, system adminstration. Sample prices:

• BEF 800/hour VAT exclusive (1/4 hour minimun) for individuals

• BEF 1500/hour VAT exclusive (One hour minimum) for corporates

Last modified:

March 15, 1997.

6.2 Phidani Software SPRL

          Rue de l'autonomie, 1
          1070 Brussels
          Belgium

Phone:

+32 2 522 06 63 Fax:

+32 2 522 09 30 EMail:

info@phidani.be URL:

http://www.phidani.be Contact:

Corchia Alain <alain@phidani.be> Type of support:

Phone, EMail, fax, remote network administration, inhouse and onsite. Special expertise:

Installation, programming services, system, administration, Internet and Intranet connectivity, firewall installation and troubleshooting.

Sample prices:

BEF 2200/hour. Last modified:

February 25, 1997.

7 Brazil

7.1 José Carlos Benfati

          Sao Paulo
          Brazil

Phone:

+55 11 8581118 EMail:

benfati@uninet.com.br Contact:

José Carlos Benfati Type of support:

Phone, EMail and inhouse. Special expertise:

Linux installation, administration and support. Intranet development and integration with existing databases and applications. Custom software development in C and C++.

Sample prices:

Upon request. Last modified:

February 19, 1997.

8 Cambodia

8.1 Mike Gärtner

          P.O. Box 177
          Phnom Penh
          Cambodia

EMail:

miky@genyosha.forum.org.kh Contact:

Mike Gärtner Type of support:

EMail and inhouse (in Cambodia, Laos, Vietnam and Thailand). Special expertise:

Internet connectivity, mail, news, UUCP. Perl. Support in English or German. Sample prices:

Upon request. Last modified:

March 20, 1997.

9 Canada

9.1 Cabala Systems Corp.

          801 3760 Albert St.
          Burnaby, BC V5C 5Y8
          Canada

Phone:

+1 604 294 1721 Fax:

+1 604 294 1751 EMail:

kcarson@direct.ca Contact:

Kevin Carson Type of support:

Support or development contracts only. Special expertise:

Administration, training and programming. We cover all Linux related areas but specialities comprise of X11 graphics, security, networking, and 24/7/365 applications.

Sample prices:

Upon request. Last modified:

March 15, 1997.

9.2 Centurion Services

          9351 Blundell Road
          Richmond, BC V6Y 1K5
          Canada

Phone:

+1 604 279 1857 Fax:

+1 604 279 1800 EMail:

censvc@ibm.net URL:

http://www.ipipeline.net/centurion Contact:

Jan Walter <jnwal@ibm.net> Type of support:

Phone, EMail, fax, remote network administration and onsite. Special expertise:

Systems design and implementation, programming. Design and construction of Linux systems, installation. Firewalling, and Internet and Intranet connectivity. Supported Linux since 1994.

Sample prices:

• CDN$75/hour for onsite labour

• CDN$55/hour for inhouse labour

• CDN$45/hour for programming (more than 15 hours)

Last modified:

April 1, 1997.

9.3 Dataden Computer Systems

          895 2nd Avenue
          Ile Bizard, QC H9C 1K3
          Canada

Phone:

+1 514 891 2293 Fax:

+1 514 696 0848 EMail:

arseneau@dataden.com URL:

http://www.dataden.com Contact:

Danny Arseneau Type of support:

Phone, EMail, fax, remote network administration inhouse and onsite.

Special expertise:

Installation, configuration, and maintenance of Linux based network servers. Installation and configuration of the Samba software package. Setting up Internet connections, firewalls, web servers, virtual web servers and most things related to Internet connections. Setting up Intranet servers. Development of CGI applications to access corporate databases. Creation of databases and web based frontends for use on Linux based web servers. We have experience in C/C++, Java, HTML, CGI, PHP/FI and Perl.

Sample prices:

CDN$60/hour. Negotiable for longer commitments. Last modified:

March 19, 1997.

9.4 Datasoft Communications

     7 Crecent Place, Suite 1822
     East York, ON M4C 5L7
     Canada

Phone:

+1 416 285 0803 Fax:

+1 416 285 5607 EMail:

mark@datasoft.on.ca URL:

http://www.datasoft.on.ca Contact:

Mark Buckaway Type of support:

Phone, EMail and inhouse. Special expertise:

Linux system configuration and installation, Internet and Intranet consulting. Please refer to our web site for additional information.

Sample prices:

CDN$85/hour (One hour minimum) or a flat rate contract. Last modified:

March 15, 1997.

9.5 Dynamicro Consulting Limited

          282 Waratah Avenue
          Newmarket, ON L3Y 5J2
          Canada

Phone:

Available to clients Fax:

Available to clients EMail:

glouis@dynamicro.on.ca Contact:

Greg Louis Type of support:

Phone, EMail and inhouse. Special expertise:

Custom software development for scientific, engineering, monitoring/control, ISP or other business applications.

Sample prices:

CDN$125/hour for short-term contracts, negotiable for longer commitments. Last modified:

March 15, 1997.

9.6 Global Proximity Corporation

          959 Alice Street
          Woodstock, ON N4S 8B1
          Canada

Phone:

+1 519 421 3541 Fax:

+1 519 421 2107 EMail:

Info@Global.Proximity.ON.CA URL:

http://Global.Proximity.ON.CA Contact:

Chris Tyler Type of support:

Full range of local and remote consulting services. Special expertise:

Advanced web and Internet development, Intranet mangement, disaster recovery preparation, and platform migrations.

Sample prices:

Base price CDN$150/hour (approximately US$120), lower for large projects, higher for urgent work. The accuracy and on-time completion of projects is guaranteed financially.

Last modified:

March 26, 1997.

9.7 G.M. Systems

     2570 Seaview Road
     Mill Bay, BC V0R 2P0
     Canada

Phone:

+1 250 743 1500 Fax:

+1 250 743 1500 EMail:

george@gmsys.com URL:

http://www.gmsys.com Contact:

George A. Farris Type of support:

Phone, EMail and inhouse. Special expertise:

Equipment sales and setup with Linux pre-installed.

Sample prices:

Upon request. Last modified:

February 20, 1997.

9.8 Hard Data Ltd.

     11060 - 166 Avenue
     Edmonton, AB T5X 1Y3
     Canada

Phone:

+1 403 456 1510 Fax:

+1 403 456 1510 EMail:

harddata@bigfoot.com Contact:

Maurice Hilarius, Michal Jaegermann Type of support:

Phone, EMail, fax, remote network administration, inhouse and onsite. Special expertise:

Installation, programming services, system administration, Internet and Intranet connectivity, firewall installation, troubleshooting, training. RAID storage and data warehousing, Samba, GIS software and support, Oracle support and connectivity, dial-up servers and remote office applications.

Sample prices:

Rates vary depending on application, level of service priority requested, and term of contract. Starting at CDN$50/hour (approximately US$40).

Last modified:

March 15, 1997.

9.9 IED Inc.

Phone:

+1 519 725 7353 Fax:

+1 519 725 7353 EMail:

orders@ied.com URL:

http://www.ied.com Contact:

Jan Vicherek Type of support:

Phone, EMail, remote network administration and inhouse. Special expertise:

Linux installation and support, network architecture design and setup (firewalls, Internet services). Intergrated web databases, online processing. Custom programming Linux and Unix, RDBMS and web.

Sample prices:

CDN$35-$60/hour, depending on type and amount of work. Last modified:

March 17, 1997.

9.10 Jones Computer Supplies

     RR \#3
     Perth, ON K7H 3C5
     Canada

Phone:

+1 613 267 6704 Fax:

+1 613 264 0372 EMail:

sos@jcs.on.ca URL:

http://www.jcs.on.ca Contact:

Peter Jones Type of support:

Phone, EMail, fax, remote network administration, inhouse and onsite. Special expertise:

Installation of Linux, network configuration (firewalls, gateways, FTP, POP and web servers). Sample prices:

CDN$60-$160/hour depending on the work being performed. Last modified:

February 20, 1997.

9.11 MostlyLinux

     Calgary, AB
     Canada

Phone:

+1 403 209 9406 Fax:

+1 403 285 1399 EMail:

info@mostlylinux.ab.ca URL:

http://www.mostlylinux.ab.ca Contact:

Scott Barker Type of support:

Phone (conditional) and EMail. Special expertise:

We do everything, from vanilla installs, to specific customizations and hacks, to ongoing system administration. However, we do not do phone-only support. We will do phone support only if we can have direct access to the system, such as through the Internet or through a no-charge dial-in. Also, since we are a very small company (at the moment), we will not return long-distance phone-calls, unless they are toll free.

Sample prices:

• $20/hour for individuals

• $40/hour for corporates Last modified:

March 15, 1997.

9.12 Gary Lawrence Murphy

     RR\#1 07 Forest Place
     Sauble Beach, ON N0H 2G0
     Canada

Phone:

+1 519 422 1150 Fax:

+1 519 422 2723 EMail:

garym@sos.on.ca URL:

http://www.sos.on.ca/˜ garym Contact:

Gary Lawrence Murphy Type of support:

Phone, EMail, remote network administration and direct or distance training (via MUD). Special expertise:

Web site installations and design, Linux installations particularly for web sites and gateways. Projects of note: Ontario Science Centre Information Highway exhibit and OSC web site, Bell Canada “AutoLinx” Intranets, SOS Internet Corp web site and ISP service, Town of Wiarton web site.

Sample prices:

CDN$75/hour but have been known to work for free or for barter for a right cause. Last modified:

March 15, 1997.

9.13 Pyx Networking

     27 Cleta Drive
     Scarborough, ON M1K 3G5
     Canada

Phone:

+1 416 264 0355 Fax:

+1 416 264 0356 EMail:

esilver@pyx.net URL:

http://www.pyx.net Contact:

E.Silver Type of support:

Phone, EMail, remote network administration, inhouse, onsite and live online. Special expertise:

Planning, design, setup, installation, maintenance as well as technical support and management of all aspects of networking using Linux alone, Linux to Novell, Unix, and Novell with Microsoft applications products. Full LAN and WAN technical support and training. Network applications development. Web systems, development. Alphanumeric paging, fax and EMail gateways. EMail listserving systems. Complete ISP network systems design, installation and maintenance. Firewalling and security systems. Installation and integration of telephone networks and CTI into new and existing computer networks. Internet and Intranet streaming audio/video broadcasting and teleconferencing systems. Specialists in Linux. Over 23 years in the computer business with a proven track record of performance. Notable local and international references available upon request.

Sample prices:

Base rates start at about US$80/hour dependent on type of work. Last modified:

March 15, 1997.

9.14 Ken Roberts Computer Consultants Inc.

     148 York Street, 2nd Floor
     London, ON N6A 1A9
     Canada

Phone:

+1 519 672 8844 Fax:

+1 519 672 3528 EMail:

krcc@mirror.org URL:

http://www.mirror.org/krcc.html Contact:

Ken Roberts Type of support:

Phone, EMail and remote network administration. Special expertise:

Custom applications software, C programming, database applications including DBase and Appgen file formats. We have worked with Unix and Linux for eight years and have must-be-reliable applications operating on systems in over 50 Ontario communities. Familiar with networks and Internet applications. Remote programming, debugging, installation and support.

Sample prices:

CDN$120/hour. Last modified:

March 15, 1997.

9.15 Silver Technologies

     Box 22123
     Regina, SK S4S 7H4
     Canada

Phone:

+1 306 596 1332 Fax:

+1 306 585 2878 EMail:

rcp@dlcwest.com Contact:

Richard Perrin Type of support:

Phone, EMail and onsite. Special expertise:

System and network administration, programming, Linux publications.

Sample prices:

CDN$20-$50/hour. Last modified:

March 27, 1997.

9.16 Troika Management Services

     Unit 611 - 1421 Western Rd
     London, ON N6G 4W4
     Canada

Phone:

+1 519 438 9323 Fax:

+1 519 438 9323 EMail:

xbill@sprynet.com URL:

http://dolphin.dbmsys.com Contact:

William McLean Type of support:

Phone, EMail, remote network administration, inhouse and onsite. Special expertise:

Job scheduling, Java, porting, system and network management, system planning, sizing and administration, installation, programming services, Internet and Intranet connectivity, firewall installation, troubleshooting.

Sample prices:

Upon request. Last modified:

April 1, 1997.

10 Denmark

10.1 Lasse Oesterild

          Noerrevang 2
          DK-4300 Holbaek
          Denmark

Phone:

+45 59451318 EMail:

lasseoe@unixzone.dk URL:

http://www.unixzone.dk Contact:

Lasse Oesterild Type of support:

Phone, EMail, fax, remote network administration, inhouse and onsite. Special expertise:

Installation, system administration, Internet and Intranet connecttibity and troubleshooting. Sample prices:

Upon request. Last modified:

March 1, 1997.

11 Europe

Companies which are not bound to a particular country within Europe.

11.1 Jochen Hayek

          Augsburger Str. 33
          D-10789 Berlin
          Germany

Phone:

+49 171 350 28 73 (mobile), +49 30 214 798 80 Fax:

+49 30 214 798.81 EMail:

Jochen Hayek@ACM.org URL:

http://www.ACM.org/˜ Jochen Hayek Contact:

Jochen Hayek Type of support:

To be negotiated - quite unrestricted. Special expertise:

Familiar with most topics Sample prices:

ECU 60/hour. Last modified:

March 16, 1997.

11.2 Tarkus AB

          Skommartorpsv. 24
          S-735 38 Surahammar
          Sweden

Phone:

+46 70 7522246 Fax:

+46 220 33907 EMail:

tarkus@tarkus.se URL:

http://www.tarkus.se Contact:

Carlos Vidal, Bo Berggren Type of support:

Phone, EMail, and inhouse if required. Special expertise:

We work in Scandinavia, Germany and Switzerland. Sample prices:

From SEK 0 to 700/hour (US$0-$100) depending on the job. Many ”easy” questions are answered for free.

Last modified:

March 15, 1997.

12 Finland

12.1 Netsol Network Solutions Oy

          Hiihtotie 3A
          FIN-01280 Vantaa
          Finland

Phone:

+358 50 5666898 Fax:

+358 9 34090404 EMail:

netsol@netsol.fi URL:

http://www.netsol.fi Contact:

Timo Virtaneva Type of support:

Inhouse. Special expertise:

Support for Internet connections, firewalls and networking. Sample prices:

Upon request. Last modified:

March 16, 1997.

13 Germany - General

13.1 Feichtinger IT Consulting

          Ziegelh\"{u}tte 4
          D-66862 Kindsbach
          Germany

Phone:

+49 6371 60056 Fax:

+49 6371 69957 EMail:

feichtinger@iccc.se Contact:

Karl Feichtinger Type of support:

Fernwartung. Special expertise:

Entwicklung und Administration von Internet und Intranet-Lösungen, Internet-Anbindung von LANs durch Linux-Kommunikationsserver, Schulung, Support.

Sample prices:

Tagessatz DM 1200, Festpreise nach Vereinbarung. Last modified:

March 16, 1997.

13.2 mdt

Phone:

+49 171 8030373 Fax:

+49 171 8029279 EMail:

mdt@contrib.de Contact:

Michael Dietrich Type of support:

Phone, EMail, inhouse. I do projects anywhere in Germany or even in other countries. Special expertise:

Programming services (various languages including C, C++ and Java), installation, troubleshooting, training. Special portable computer knowledge.

Sample prices:

Maximum of DM 100/hour (approximately US$/60) plus traveling. Last modified:

March 23, 1997.

13.3 MetaWorks GmbH

          Carl-Sp\"{a}ter-Str. 1
          D-56070 Koblenz
          Germany

Phone:

+49 261 98440 0 Fax:

+49 261 98440 44 EMail:

support@metaworks.de URL:

http://www.metaworks.de Contact:

Andreas Heilwagen Type of support:

Phone, EMail and inhouse. Special expertise:

Hostmaster of an ISP using Linux machines. Sample prices:

Upon request. Last modified:

March 19, 1997.

13.4 Ntakpe Management & Consulting

          Waldweg 9
          D-85386 Eching/Munich
          Germany

Phone:

+49 89 327155 0 Fax:

+49 89 327155 99 EMail:

ntakpe@ntakpe.com URL:

http://www.ntakpe.com Contact:

Jean Louis Ntakpe Type of support:

Phone, EMail and inhouse. Special expertise:

Familiar with most topics. Sample prices:

Upon request. Last modified:

March 17, 1997.

13.5 S.u.S.E. Gmbh

          Gebhardtstr. 2.
          D-90762 F\"{u}rth
          Germany

Phone:

+49 911 32471 56, +49 911 7405331 Fax:

+49 911 7405331 EMail:

psupport@suse.de URL:

http://www.suse.de Contact:

Klaus Franken, Bodo Bauer Type of support:

Telephonisch, EMail, Fax, Fernwartung, Inhouse und vor Ort. Special expertise:

Support mit/ohne Wartungsverträge, Entwicklung, WWW-Applikationen, Internet- und Intranet-Anbindung (z.B: ISDN), spezielle Hardwareanbindung, Beratung. Umfassendes Linux Know-How des 12-köpfigen Entwicklerteams der S.u.S.E. Linux Distribution.

Sample prices:

• DM 1.500 für die Installation Linux-Rechner als Internetserver (WWW, EMail, news, proxy, Paketfilter-Firewall)

• DM 140/Monat für Support/Wartungsvertrag eines Intranet-Servers

• DM 1.500/Manntag für Installation/Beratung

Weitere Beispiele finden Sie unter http://www.suse.de/Supportuebersicht.html Last modified:

March 27, 1997.

14 Germany - Baden-Württemberg

14.1 Hans-Joachim Baader

     Poppelfeld 9
     D-76646 Bruchsal
     Germany

Phone:

+49 7257 930142 Fax:

+49 7257 930143 EMail:

hans@grumbeer.inka.de Contact:

Hans-Joachim Baader Type of support:

Free support by EMail, phone, fax if I can answer by EMail. Special expertise:

Linux and Unix programming and system administration. Process automation, network programming, databases, realtime systems.

Sample prices:

Upon request. Last modified:

March 15, 1997.

14.2 ISS Internet Service Stuttgart GmbH

     Wilhelm Schickard-Str. 7
     D-72124 Pliezhausen
     Germany

Phone:

+49 7127 9740 Fax:

+49 7127 974110 EMail:

info@neckar-alb.de URL:

http://www.neckar-alb.de Contact:

Jochen Wiedmann <Jochen.Wiedmann@Neckar-Alb.DE> Type of support:

Phone, EMail, remote network administration, inhouse and onsite. Special expertise:

Firewall installation, programming (C, C++, Perl), system administration, Internet, Intranet and training.

Sample prices:

DM 160/hour. Last modified:

March 19, 1997.

14.3 MU Softwareentwicklung

          Wacholderweg 14
          D-70597 Stuttgart
          Germany

Phone:

+49 711 7676184 Fax:

+49 711 7676185 EMail:

info@musoftware.com URL:

http://www.musoftware.com (In German) Contact:

Ullrich von Bassewitz <uz@musoftware.com> Type of support:

Phone, EMail and inhouse. Special expertise:

Installation, programming services, Internet and Intranet connectivity. Sample prices:

Upon request. Last modified:

March 16, 1997.

14.4 Jochen Wiedmann

          Am Eisteich 9
          D-72555 Metzingen
          Germany

Phone:

+49 7123 14881 Fax:

+49 7127 974110 EMail:

Jochen.Wiedmann@Neckar-Alb.DE Contact:

Jochen Wiedmann Type of support:

Phone, EMail, remote network administration and inhouse. Special expertise:

Firewall installation, programming (C, C++, Perl), system administration, network, Internet, Intranet.

Sample prices:

DM 120/hour. Last modified:

March 19, 1997.

14.5 Michael Will Linux Systeme

     Fichtenweg 1/1
     Room 312
     D-72076 T\"{u}bingen
     Germany

Phone:

+49 7071 969063 Fax:

+49 7071 969064 EMail:

Michael.Will@student.uni-tuebingen.de URL:

http://www-ti.uni-tuebingen.de/˜ will Contact:

Michael Will Type of support:

Phone, EMail, remote network administration and inhouse. Special expertise:

Installation, programming services, system administration, Internet and Intranet connectivity, firewall installation, troubleshooting and training.

Sample prices:

Individual pricing for the commercial market, feel free to ask. In general, I do free support for home users and universities with light weight troubleshooting and consulting. When I have to invest more time, I charge money.

Last modified:

March 16, 1997.

15 Germany - Bayern

15.1 B.I.S Babylon Internet Service GmbH

     Runde-Turm-Str. 2a
     D-63785 Obernburg
     Germany

Phone:

+49 6022 6160 20 Fax:

+49 6022 6160 16 EMail:

babylon@babylon.de URL:

http://www.babylon.de Contact:

Andreas Tikart Type of support:

Schulungen in unserem Schulungsraum mit Internet-Standleitung, u.a. zu Linux und Aufbau eines Dial-In Servers mit Linux.

Special expertise:

Wir betreiben einen Internet-Einwählknoten (POP) unter Linux. Sample prices:

Auf Anfrage. Last modified:

March 25, 1997.

15.2 EDV-Beratung Gert Döring

          Frankfurter Ring 24
          D-80807 M\"{u}nchen
          Germany

Phone:

+49 89 3545979 Fax:

+49 89 3545980 EMail:

gert@greenie.net Contact:

Gert Döring, Simone Demmel Type of support:

remote network administration, inhouse and onsite. Special expertise:

Networking, LAN/WAN planning and implementation, system integration, system programming, training, general support.

Sample prices:

Prices vary depending on the job. Prices can be negotiated.

• Low prices for non-profit organizations

• DM 95/hour for short jobs

• DM 700/day for longer jobs

Last modified:

March 16, 1997.

15.3 K&K Kommunikationssysteme GmbH

          Haunstetterstr.19
          D-86161 Augsburg
          Germany

Phone:

+49 821 25892 20 Fax:

+49 821 25892 22 EMail:

Alexander.Mueller@Exklusiv.De URL:

http://www.exklusiv.de Contact:

Alexander Müller, Oliver Kaufmann, Felix Kirschner Type of support:

Remote network administration and onsite. Special expertise:

Installation, Internet connectivity, Perl, shell, networking, ISDN with Linux. Sample prices:

DM 150/hour. Last modified:

March 17, 1997.

15.4 noris network GmbH

          Schleiermacherstra{\ss}e 12
          D-90491 N\"{u}rnberg
          Germany

Phone:

+49 911 59818 0 Fax:

+49 911 59818 11 EMail:

info@noris.de URL:

http://www.noris.de Contact:

Matthias Urlichs <smurf@noris.de> Type of support:

Phone, EMail, inhouse and remote network administration.

Special expertise:

Extensive kernel, debugging and networking experience. Specializes on system-level drivers, tools and utilities.

Sample prices:

DM 150/hour for phone or inhouse support. Last modified:

March 15, 1997.

15.5 Julian Howard Stacey, Vector Systems Limited, München

     Holz Stra{\ss}e 27d
     D-80469 Munich
     Germany

Phone:

+49 89 268616 Fax:

+49 89 2608126 ISDN Data:

+49 89 26023276 EMail:

jhs@freebsd.org URL:

http://www.freebsd.org/˜ jhs/vsl Contact:

Julian Stacey Type of support:

Phone (only after prepaid contract), EMail Special expertise:

Training at customer site custom development on and offsite, system config, system administration, etc. Mainly FreeBSD based, but also Linux. I provide consulting in Bayern (Germany) and South East England.

Sample prices:

Upon request. Last modified:

March 16, 1997.

15.6 WorNet Internetdienste

          Weidacher Hauptstr. 61
          D-82515 Wolfratshausen
          Germany

Phone:

+49 8171 22937 Fax:

+49 8171 22938 EMail:

wornet@wor.de URL:

http://www.wor.de Contact:

Christian Eich Type of support:

Phone, EMail and inhouse. Special expertise:

Internet and Intranet server, ISDN4Linux, dial-in server, file and print server (Unix and Windows), web applications.

Sample prices:

DM 80/hour. Last modified:

March 3, 1997.

16 Germany - Berlin

16.1 netcomp Informationsmanagement

          Chausseestr. 85
          D-10115 Berlin
          Germany

Phone:

+49 30 28390576, +49 171 2710881 (mobile)

Fax:

+49 30 28390577 EMail:

fruehbrodt@msn.com, 100115.576@compuserve.com Contact:

Jörg Frühbrodt Type of support:

Remote network administration, inhouse and onsite. Special expertise:

Beratung, Entwicklung, Vertrieb, Installation, Service und Support für Linux; auch Komplettlösungen.

Sample prices:

• DM 120/Stunde für onsite Support, Tages- und Wochen-Preise nach Vereinbarung • DM 220/Monat mit Support-Vertrag

Last modified:

April 9, 1997.

16.2 Woller & Link GbR

          Grunewaldstra{\ss}e 9
          D-10823 Berlin
          Germany

Phone:

+49 3009 98 53, +49 172 3124656 (mobile) Fax:

+49 30 21750288 EMail:

wl@inx.de URL:

http://www.inx.de/w+l Contact:

Daniel Link Type of support:

Telephonisch, EMail, Fax und Fernwartung. Special expertise:

System-Administration und Beratung, Netzwerkplanung und Aufbau, Vermittlung von Internetzugängen, Verkauf von für Linux geeigneten Hardwarekomponenten und Systemen.

Sample prices:

Informationen über die aktuelle Preisgestaltung liegen auf unserer Website. Last modified:

March 16, 1997.

17 Germany - Brandenburg

17.1 Bernd Hentig EDV-Beratung

     Parkstr. 10
     D-16244 Finowfurt
     Germany

Phone:

+49 3335 32330 Fax:

+49 3335 32331 EMail:

Bernd.Hentig@guug.de Contact:

Bernd Hentig Type of support:

Phone, EMail, fax and inhouse. Special expertise:

LAN/WAN planning and installation (Ethernet, ISDN, X.25), software development and integration of standard software, Internet and Intranet connectivity and presentation. Sale and support of preinstalled Linux hardware (Intel). Sale and installation of standard software components (e.g. Red Hat Linux, ApplixWare).

Sample prices:

• DM 50/hour for assembling web pages (including scanning)

• DM 80/hour for Unix and Linux development jobs

• DM 120/hour for installation and support jobs

• DM 160/hour if non-Unix systems are involved (Novell, Windows)

• Special prices for larger projects on request

Last modified:

March 19, 1997.

18 Germany - Hamburg

18.1 Jens Doll

          Hauke-Haien-Weg 29
          D-Hamburg 22149
          Germany

Phone:

+49 40 67391235 Fax:

+49 40 67391236 EMail:

100117.3615@compuserve.com URL:

http://ourworld.compuserve.com/homepages/renate Contact:

Jens Doll Type of support:

Inhouse and training. Special expertise:

Application development, networking. Sample prices:

Upon request. Last modified:

March 25, 1997.

18.2 Dipl.-Ing. Axel Ramge

          Ebertallee 16
          D-22607 Hamburg
          Germany

Phone:

+49 40 893060 Fax:

+49 40 893061 EMail:

axel@ramge.de URL:

http://www.ramge.de/ax Contact:

Axel Ramge Type of support:

Telephonisch. Special expertise:

Alles, mit Schwerpunkt Ingenieurbüro: Analyse, Auswahl, Beschaffung, Installation, Wartung und Betreuung. Software: CAD, Datenbanken, Technische Software, Kommunikation (Mail, Fax, WWW).

Sample prices:

• DM 10/Telephonat (Hotline)

• DM 110/hour

• Tages- und Projektsatz nach Vereinbarung

Last modified:

March 20, 1997.

19 Germany - Hessen

19.1 Gerd Aschemann

          Osannstr. 49
          D-64285 Darmstadt
          Germany

Phone:

+49 6151 16 2259 Fax:

+49 6151 16 3052 EMail:

aschemann@Informatik.TH-Darmstadt.de URL:

http://www.informatik.th-darmstadt.de/˜ ascheman Contact:

Gerd Aschemann Type of support:

Phone, EMail and inhouse. Sample prices:

US$100/hour. Last modified:

February 19, 1997.

19.2 b.i.t. beratungsgesellschaft für informations-technologie mbh

          Bunsenstrai{\ss}e 15
          D-64293 Darmstadt
          Germany

Phone:

+49 6151 827575 Fax:

+49 6151 827576 EMail:

info@b-i-t.de URL:

http://www.b-i-t.de Contact:

Reinhold Schucht, Andy Schmidt Type of support:

Telephonisch, EMail und vor Ort. Special expertise:

Beratung bei Planung und Durchführung von Softwareprojekten, Projektleitung,

Softwareentwicklung, Administration von Unix-Netzwerken und Internet-Anbindung, Schulung. Sample prices:

DM 120/hour. Last modified:

March 20, 1997.

19.3 LEMIS

          Schellnhausen 2
          D-36325 Feldatal
          Germany

Phone:

+49 6637 919123 Fax:

+49 6637 919122 EMail:

grog@lemis.de URL:

http://www.lemis.de/˜ grog Contact:

Greg Lehey Type of support:

Phone, EMail, fax and inhouse. Special expertise:

Familiar with most topics. Sample prices:

Typically US$100 or DM 150/hour. Significant reductions for non-profit organizations. Last modified:

February 20, 1997.

20 Germany - Mecklenburg-Vorpommern

20.1 Bernd Wolf

     M\"{o}llner Str. 11, Wh. 3, Zi. 45a,
     D-18109 Rostock
     Germany

Phone:

+49 177 2350472 EMail:

wolf@egd.igd.fhg.de URL:

http://www.egd.igd.fhg.de/˜ wolf Contact:

Bernd Wolf Type of support:

Phone and EMail. Special expertise:

Verschiedenes, z.B.: System Installation und Administration, ”troubleshooting”, Hard-, Softwareund Netz-Tips, Hilfe: learning by doing, außerhalb Rostock: kommerziell (bis auf EMail) - nach Absprache, größere Aufträge.

Sample prices:

Auf Anfrage. Last modified:

February 19, 1997.

21 Germany - Niedersachsen

21.1 Gärtner Datensysteme

     Hamburger Str. 273a
     D-38114 Braunschweig
     Germany

Phone:

+49 531 2335555 Fax:

+49 531 2335556 EMail:

info@gaertner.de URL:

http://www.gaertner.de Contact:

Joachim Beggerow, Martin Neitzel Type of support:

Phone, EMail and inhouse. Special expertise:

Familiar with most topics. Sample prices:

Depending on the job, starting at DM 1800 for a full day. Last modified:

March 16, 1997.

21.2 Jürgen Gerloff Elektronik

          Th\"{o}rener Str. 14
          D-29308 Winsen
          Germany

Phone:

+49 5146 92046 Fax:

+49 5146 92047 EMail:

postmaster@gerloff.de Contact:

Jürgen Gerloff Type of support:

Telephonisch, Fernwartung und vor Ort.

Special expertise:

Verkauf, Installation, Service und Administration von Linux-Systemen. Sample prices:

Auf Anfrage. Last modified:

March 16, 1997.

21.3 Electronic & Computertechnik Margitta Holzbrecher

          M\"{u}hlenkampstr. 24
          D-31515 Wunstorf
          Germany

Phone:

+49 5031 12464 Fax:

+49 5031 12464 EMail:

M.Holzbrecher@t-online.de URL:

http://home.t-online.de/home/M.Holzbrecher Contact:

Margitta Holzbrecher Type of support:

Phone, EMail and inhouse. Special expertise:

Linux consulting, installation , administration and Linux PCs.

Linux Beratung, Installation, Administration und Komplettsysteme. Sample prices:

DM 80/hour. Last modified:

March 17, 1997.

21.4 Martin Schulze

          Infodrom Oldenburg
          Artillerieweg 55a
          D-26129 Oldenburg
          Germany

Phone:

+49 441 7780584 Fax:

+49 441 777884 EMail:

joey@infodrom.org URL:

http://home.pages.de/˜ joey Contact:

Martin Schulze Type of support:

Phone, EMail and inhouse. Special expertise:

Familiar with most topics. Sample prices:

Upon request. Last modified:

March 15, 1997.

22 Germany - Nordrhein-Westfalen

22.1 Stephan Austermühle

          Gl\"{u}ckaufstr. 52/053
          D-57076 Siegen
          Germany

Phone:

+49 271 790504 Fax:

+49 271 790532 EMail:

sa@netaccess.de URL:

http://www.wad.org/users/stephan Contact:

Stephan Austermühle Type of support:

Phone (only if service contract exists), EMail and inhouse. Special expertise:

I serve the regions Kreis Altenkirchen, Cologne, Oberbergischer Kreis, Rhein-Sieg-Kreis, Siegerland, and Westerwaldkreis.

Sample prices:

• DM 80/hour for hotline support

• DM 5,000/year for a maintenance contract

• DM 1,500 (typically) for installing a Linux-based Internet or Intranet server

Last modified:

March 19, 1997.

22.2 dev/consulting GmbH

          Webereistr. 3
          D-33602 Bielefeld
          Germany

Phone:

+49 521 1365800 Fax:

+49 521 1365803 EMail:

devcon@devconsult.de URL:

http://www.devconsult.de Contact:

Winfried Motzkus, Christian Carstensen Type of support:

Phone, EMail, remote network administration and inhouse. Special expertise:

Installation, administration, Internet administration, object oriented database software engeneering and support of custom solutions on the base of Linux. Security check, development of Internet and Intranet solutions.

Sample prices:

• DM 110/hour

• Support on monthly rates of DM 250/month available

• Development on fixed prices rate

Last modified:

March 31, 1997.

22.3 Ing. Büro Gerichhausen

          Steinsstra{\ss}e 103
          D-41199 M\"{o}nchengladbach
          Germany

Phone:

+49 2166 15068 Fax:

+49 2166 10159 EMail:

support@mail.niederrhein.de URL:

http://www.niederrhein.de Contact:

Christoph Adomeit Type of support:

Phone, EMail, remote network administration and inhouse. Special expertise:

Internet, firewalls, ISDN, TCP/IP, proxy, mail and news, UUCP and security.

Sample prices:

DM 155/hour. Last modified:

February 19, 1997.

22.4 netServe oHG

     Michaelweg 66
     D-48149 M\"{u}nster
     Germany

Phone:

+49 251 866 976 Fax:

+49 251 866 977 EMail:

info@www.muenster.com URL:

http://www.muenster.com Contact:

Jörg Fraune Type of support:

Telephonisch, EMail, Fax und Fernwartung. Special expertise:

Systemverwaltung, LAN und WAN, Sicherheitschecks, Linux-Schulungen, Softwareentwicklung, Firewalls, heterogene Netzwerke mit Linux-Server und Windows-Clients, Anwendungsentwicklung, Netzwerkinstallation und -wartung im Rahmen von Schulen im Netz, WWW-Design, Webserverinstallation, ISDN-Intranet Realisierung.

Sample prices:

• Stundensatz DM 120

• Linux Komplettinstallation: DM 600

Nähere Preisinformationen auf Anfrage und auf unserer Web Server. Last modified:

April 8, 1997.

22.5 PiN - Präsenz im Netz GmbH

          Leimbacher Str. 36
          D-42281 Wuppertal
          Germany

Phone:

+49 202 2 50 11 64 Fax:

+49 202 2 50 11 65 EMail:

pin@pinserve.de URL:

http://www.pinserve.de Contact:

Kurt Harders, Andre Dressler Type of support:

Inhouse. Special expertise:

Training, site planning, Intranet and Internet, client/server, database integration, application development, GNU Software.

Sample prices:

Upon reqest. Last modified:

March 15, 1997.

22.6 Silent Running

          Hospitalstra{\ss}e 2
          D-53840 Troisdorf
          Germany

Phone:

+49 2241 87190 Fax:

+49 2241 871946 EMail:

info@silent-running.de Contact:

Ralf Nitzgen, Ray Pye Type of support:

Phone, EMail, fax, remote network administration, inhouse and onsite. Special expertise:

Specializing in object-oriented software engeneering and networking under Linux. Services include design, development and support of custom systems (particularly business systems), the provision of low-cost Internet and Intranet connectivity solutions (firewalls and gateways) and general Linux support (installation, configuration, security and tuning).

Sample prices:

DM 115/hour. Fixed price basis available for development projects. Last modified:

March 19, 1997.

22.7 SUNSOFT Büro und EDV GmbH

     M\"{u}hlenbach 42
     D-50676 K\"{o}ln
     Germany

Phone:

+49 221 9230580 Fax:

+49 221 9230581 EMail:

buero@sunsoft.de URL:

http://www.sunsoft.de Contact:

Egbert Hinzen Type of support:

Phone, EMail, remote network administration and inhouse.

Special expertise:

Installation, system administration, Internet and Intranet connectivity, firewall installation, troubleshooting and training.

Sample prices:

DM 100/hour. Last modified:

April 6, 1997.

22.8 transdata - Bigge & Langanke GbR

          Laventiestr. 17
          D-58640 Iserlohn
          Germany

Phone:

+49 2371 963616 Fax:

+49 2371 461381 EMail:

bigge@transdata.de URL:

http://transdata.de Contact:

Frank Bigge, Boris Langanke Type of support:

EMail and inhouse. Special expertise:

Training, installation and administration, Internet administration, web authoring, web design and 3D modelling.

Sample prices:

• DM 120/hour or DM 1200/day for installation and administration

• DM 120/hour for training

• DM 150/hour for web authoring and web design

Last modified:

March 20, 1997.

22.9 Jörg Weber

          Sursstr.1
          D-45143 Essen
          Germany

Phone:

+49 172 2186328 Fax:

+49 201 312677 EMail:

kufi@kufis.ruhr.de Contact:

Jörg Weber Type of support:

Phone, EMail and remote network administration. Special expertise:

Training, system installation and maintainance (Standalone, LAN, WAN). Troubleshooting. Sample prices:

• Free for privates and non-profit organizations if it doesn’t take too much time

• In general DM 120/hour plus expenses and VAT

Last modified:

February 19, 1997.

23 Germany - Saarland

23.1 B. Neubert EDV-Dienstleistungen

          Parkweg 16
          Scheidt
          D-66133 Saarbr\"{u}cken
          Germany

Phone:

+49 681 817629 Fax:

Upon request.

EMail:

neubert@lusi.uni-sb.de (nur für private nichtkommerzielle und wissenschaftliche Kontakte, for private non-commercial and scientific contacts only)

Contact:

Boris Neubert Type of support:

Telephonisch, Inhouse und vor Ort. Special expertise (German):

Die Beratungstätigkeit umfaßt Hilfestellung bei Hardware- und Software-Kauf sowie Installation von Hardwarekomponenten und Anwenderprogrammen, Netzwerklösungen im Unix-Bereich, Vermittlung von Netzzugängen, sowie die allgemeine Informationsbeschaffung im wissenschaftlichen Bereich und verwandte Dienstleistungen.

Special expertise (English):

Consulting comprises assistance with hardware and software purchase, installation of hardware components and user programs, networking solutions under Unix, mediation of network access, general information retrieval in the scientific field and related services.

Sample prices:

Auf Anfrage; upon request. Last modified:

March 17, 1997.

24 Germany - Schleswig-Holstein

24.1 AllCon GmbH

     Lise Meitner Stra{\ss}e 2
     D-24941 Flensburg
     Germany

Phone:

+49 461 9992 162 Fax:

+49 461 9992 165 EMail:

info@allcon.net URL:

http://www.allcon.net Contact:

Oliver J. Albrecht, Göran Jensen, Dieter Thiel Type of support:

Phone, remote network administration and inhouse. Special expertise:

Installation and system administration, router and firewall installation, setup and maintenance of Linux and Unix based database systems and web servers, Internet and Intranet services and connectivity, training.

Sample prices:

DM 120/hour. Last modified:

March 16, 1997.

25 Germany - Thüringen

25.1 IKS GmbH

     Wildenbruchstra{\ss}e 15
     D-07745 Jena
     Germany

Phone:

+49 3641 675640 Fax:

+49 3641 675655 EMail:

info@iks-jena.de URL:

http://www.iks-jena.de Contact:

Erik Heinz Type of support:

Phone, EMail and remote network administration.

Special expertise:

Installation, training. We also do network planning and setup, programming, and provide full Internet service.

Sample prices:

DM 150/hour for consulting. Contract prices available. Last modified:

March 17, 1997.

26 Hungary

26.1 Krisztián Czakó

     52 Fodor Street
     H-1124 Budapest
     Hungary

Phone:

+36 1 2144070 Fax:

+36 1 1566088 EMail:

slapic@fido.hu Contact:

Krisztián Czakó Type of support:

Phone, EMail, fax, remote network administration, inhouse and onsite. Special expertise:

Installation of Internet and Intranet servers (e.g. web, FTP, mail, samba, mars), office systems (based on Applixware and maybe on Caldera’s products in the near future). Of course I can install many of the free software packages as well. Training.

Sample prices:

Upon request. Last modified:

March 6, 1997.

27 India

27.1 Algologic Research & Solutions

          78 Sancharpuri Colony
          Bowenpally P.O.
          Secunderabad 500 011
          India

Phone:

+91 40 775 1650 Fax:

+91 40 842 477 (Please mention ”Attention FMB 229”) EMail:

algolog@hd1.vsnl.net.in Contact:

S. Parthasarathy Type of support:

Phone, EMail, onsite within India and neighbourhood countries. Special expertise:

Familiar with most topics. Sample prices:

Upon request. Last modified:

March 22, 1997.

28 Israel

28.1 TclTek Ltd.

          Mehalke HaMayim 12
          Jerusalem 91112
          Israel

Phone:

+972 52 670 353 Fax:

+972 2 563 0948 EMail:

benavrhm@tcltek.co.il URL:

http://www.tcltek.co.il Contact:

Jonathan Ben-Avraham Type of support:

Phone, EMail and onsite in Israel, Jordan and Palestinian Autonomous areas. Support provided in English, Hebrew and some Arabic.

Special expertise:

Custom firewalls and routers, secure servers. Sample prices:

US$50/hour for companies, free via linux-il@linux.org.il Last modified:

March 16, 1997.

29 Italy

29.1 P.O.P. Art S.r.l.

          Via Camperio, 9
          I-Milano
          Italy

Phone:

+39 338 8164150 Fax:

+39 2 57605058 EMail:

multivac@comm2000.it Contact:

Luca Pescatore Type of support:

EMail, remote network administratiom, inhouse and onsite.

Special expertise:

PC (Intel/PowerPC) with Linux (Hardware), system administration, Internet and Intranet Connectivity, web administration and Installation and firewall configuration.

Sample prices:

Upon reqeust. Last modified:

January 12, 1997.

29.2 PROFUSO

     Via Vergani, 11
     I-35031 Abano Terme (Padova)
     Italy

Phone:

+39 348 2220811 EMail:

profuso@profuso.com URL:

http://www.profuso.com Contact:

Giuseppe Zanetti <beppe@profuso.com> Type of support:

Phone, EMail, onsite only in the province of Padova and in Veneto. Special expertise:

Located in Veneto. Support both Intel and Alpha Linux systems. System Integration and LAN technologies. Intranet. We offer Linux installation and Linux based solutions (for example ISDN routers). We are releasing a very important product. Please see our web site for more info.

Sample prices:

Upon request. Last modified:

March 19, 1997.

29.3 SIR s.r.l. - PublinetWork

          Piazza Tavarone 2/22
          I-16123 Genova
          Italy

Phone:

+39 10 2463638 Fax:

+39 10 2468561 EMail:

info@publinetwork.com URL:

http://www.publinetwork.com Contact:

Luca Maranzano Type of support:

Phone, EMail, remote network administration and inhouse (when possible). Special expertise:

Training, Internet and Intranet connectivity, web publishing, ready to run Linux servers with installation and configuration.

Sample prices:

Upon request. Last modified:

March 19, 1997.

29.4 Softmedia Spa

          via Pietrasanta 14
          I-20141 Milano
          Italy

Phone:

+39 335 417677 Fax:

+39 2 5730 7652 EMail:

venta@softmedia.it URL:

http://www.softmedia.it Contact:

Paolo Ventafridda Type of support:

Phone, remote network administration and inhouse. Special expertise:

We offer a custom preinstalled Linux box with some special patches, to be installed on LANs for caching proxy, IP Masquerading, DNS, SMTP, POP mail and dial on demand, using ethernet routers such as LanOptics Ascend, or modems, or Teles ISDN cards. It’s a plug and play solution we sell as a service, not as a product. The Linux box comes with a triple DES filesystem encryption. We help selecting the best ISP, we negotiate the best prices, for quality cost/effective connections. We do system programming, with 12 years experience in C/Unix . We are focused on Intranet solutions, with an eye on Linux integration with Mac and Windows 95/NT. We can and do provide serial installation for many Linux boxes, and assist our customers with training on security, system administration and application programming. Finally we integrate Linux platforms with NT fax on demand software, and voice processing systems based on Dialogic cards.

Sample prices:

Upon request. Last modified:

March 9, 1997.

29.5 Studio LEADER Pro

     Via Pietrastretta, 76
     I-38100 Trento
     Italy

Phone:

+39 461 828229 Fax:

+39 461 829862, +39 461 829877 EMail:

info@leader.it URL:

http://www.infomagic.com/leader.it, http://www.leader.it Contact:

Guido Brugnara <brugnara@leader.it>, Isabella Faggioni <faggioni@leader.it> Type of support:

EMail, remote network administration and inhouse. Special expertise:

Installation, engineering. LAN and WAN integration. Support for InfoMagic clients. Sample prices:

Upon request. Last modified:

April 11, 1997.

30 Japan

30.1 SoftWare Research Co.

          Hayabusacho 2-11
          Chiyodaku, Tokyo 102
          Japan

Phone:

+81 3 3264 6171 Fax:

+81 3 3264 0647 EMail:

info@src.co.jp URL:

http://www.src.co.jp Contact:

Hayama Yossihiro Type of support:

EMail.

Special expertise:

Linux installation and configuration, Internet and Intranet conectivity. Sample prices:

Upon request. Last modified:

March 17, 1997.

31 Mexico

31.1 Miguel de Icaza

          Av Copilco 163-22-1003
          Mexico City
          Mexico

Phone:

+52 915 622 4680 EMail:

miguel.de.icaza@linux.org Contact:

Miguel de Icaza Type of support:

Phone and EMail. Special expertise:

Kernel hacking, driver developement, system administration, applications programming, bug fixing of publicaly available software, extension of freely redistributable software.

Sample prices:

US$30-$50/hour. Last modified:

February 19, 1997.

31.2 Laboratorio de Sistemas Linux

          Calzada de los Gallos \#20
          Edificio 8 Departamento 6
          Fraccionamiento Villas de Cortes
          Tejalpa, Morelos 62550
          Mexico

Phone:

+52 73 170952 Fax:

+52 73 297098 EMail:

turbo@linux.lsl.com.mx URL:

http://www.lsl.com.mx Contact:

Aztec Eagle Turbo, IRL Hector Daniel Cortes Gonzalez Type of support:

EMail and onsite. Special expertise:

General Linux consulting, Linux workstations, networking, Internet connectivity, and web server. Sample prices:

Upon request. Last modified:

March 16, 1997.

32 The Netherlands

32.1 Berrymount Automation B.V.

          P.O. Box 3362
          NL-2601 DJ Delft
          The Netherlands

Phone:

+31 15 2145985 Fax:

+31 15 2122179 EMail:

info@berrymount.nl URL:

http://www.berrymount.nl Contact:

Rob van Nieuwkerk Special expertise:

Consultancy, projects, services and development in technical automation, data- and telecom, high performance computing and computer security. Berrymount has strong Linux expertise.

Sample prices:

Upon request. Last modified:

March 16, 1997.

32.2 Bitbybit Information Systems

          Kluyverweg 2A
          NL-2629 HT Delft
          The Netherlands

Phone:

+31 15 2682569 Fax:

+31 15 2682530 EMail:

info@bitbybit-is.nl URL:

http://www.bitbybit-is.nl Contact:

Erno Marks Type of support:

Phone, EMail, remote network administration and onsite.

Special expertise:

Installation, troubleshooting, training, networking, database systems, timetabling systems. Sample prices:

US$60-$75/hour. Last modified:

March 17, 1997.

32.3 BitWizard B.V.

          van Bronckhorststraat 12
          NL-2612 XV Delft
          The Netherlands

Phone:

+31 70 3700841 Fax:

+31 15 2138217 EMail:

info@BitWizard.nl URL:

http://www.BitWizard.nl Contact:

Rogier Wolff <R.E.Wolff@BitWizard.nl> Type of support:

Phone, EMail, remote network administration, inhouse and onsite. Special expertise:

Troubleshooting, programming, firewall installation, Linux installation, device drivers. Sample prices:

• US$80/hour, but we usually make you a fixed price offer before we start work.

• US$3200-$6400 for a device driver.

Last modified:

March 17, 1997.

32.4 Cistron Internet Services

          P.O. Box 297
          NL-2400 AG
          The Netherlands

Phone:

+31 172 419445 Fax:

+31 172 430979 EMail:

linux@cistron.nl URL:

http://www.cistron.nl Contact:

Danny ter Haar Type of support:

Phone, EMail, fax, remote network administration, inhouse and onsite. Special expertise:

Author of large parts of Linux (sysvinit, minicom, network code and kernel code). Maintainer of many Debian GNU/Linux packages. Internet and Intranet provider including firewalls and leased lines.

Sample prices:

HFL 160/hour. Last modified:

February 22, 1997.

32.5 MANTRA Research

          Wijenburg 10
          NL-5655 EW Eindhoven
          The Netherlands

Phone:

+31 40 2910125 Fax:

+31 40 2910126 EMail:

sales@mantra.nl URL:

http://www.mantra.nl Contact:

Naresh Sharma Type of support:

Consultancy on per hour basis. Maintenance contracts for remote system administration. Turn-key installations. Sales of Caldera, and related software. Inhouse development of software solutions (we do not do datacheering).

Special expertise:

Design automation, engineering design and analysis, turn-key installations, firewalls, cryptography, networking, development. We install Linux on all types of processor architectures, multiprocessor Intel machines, SPARCs and Alphas. We are a Caldera Channel Partner.

Sample prices:

Rates are on per hour basis, they exclude BTW (VAT) and the transportation charges for a consultant, fixed prices by negotiation. These prices are an indication and not binding. For maintenance contracts, please send us an EMail or phone us. Site maintenance is provided for installations with more than 100 workstations. The first hour of consultancy for any new client is free.

• Hfl 125/hour (approximately US$70) for Linux consulting

• Hfl 195 (approximately US$100) for cryptography, firewalls

• Hfl 175-250 (approximately US$90-$135) for engineering design (FEA, modelling, analysis)

• Hfl 395 (US$200) for design automation

Last modified:

March 16, 1997.

32.6 Netco

     Zijlstraat 78/80
     NL-2011 TR Haarlem
     The Netherlands

Phone:

+31 23 5345554 Fax:

+31 23 5316155 EMail:

info@netco.nl URL:

http://www.netco.nl Contact:

Martijn van der Ende Type Type of support:

Phone, EMail, remote network administration, inhouse and onsite. Special expertise:

Internet and Intranet connectivity, firewall installation. Sample prices:

Hfl 125/hour. Project are done on fixed price. Last modified:

April 8, 1997.

32.7 Le Reseau netwerksystemen BV

          Bieslookstraat 31
          NL-9731 HH Groningen
          The Netherlands

Phone:

+31 50 5492701 Fax:

+31 50 5492310 EMail:

info@reseau.nl URL:

http://www.reseau.nl Contact:

Arthur Donkers <arthur@reseau.nl> Type of support:

Phone, EMail and onsite.

Special expertise:

Fault tolerant Linux systems, high reliability Intranet and Internet servers, application development and migration, security.

Sample prices:

Dfl 115/hour, projects are done on a fixed price basis. Last modified:

April 4, 1997.

33 Norway

33.1 it-analyse as

          Breist{\o}lsveien 14
          N-5035 Bergen-Sandviken
          Norway

Phone:

+47 55962221 Fax:

+47 55962858 EMail:

magnus@login.eunet.no URL:

http://login.eunet.no/˜ magnus Contact:

Magnus Alvestad Type of support:

Phone, EMail and inhouse. Special expertise:

Emacs, Java, GNU tools, GCC, kernel, Internet, SGML. Sample prices:

US$80-$100/hour or free (if interesting work). Last modified:

March 16, 1997.

33.2 NorConnect Internet Services AS/NorConnect Group

     Boks 9273 Gr\"{o}nland
     N-0134 Oslo
     Norway

Phone:

+47 22 17 41 70 Fax:

+47 22 17 64 99 EMail:

post@norconnect.no URL:

http://www.norconnect.no/english Contact:

Vidar Hokstad <vidarh@norconnect.no> Type of support:

Phone, EMail and inhouse. Special expertise:

NorConnect specialize in planning, implementation and maintenance of Internet and Intranet related networks and server solutions. Among our reference customers you’ll find for instance IDG New Media - International Data Groups Norwegian ISP - for which we run most of the network and news, mail and authentication servers.

Sample prices:

• Starting at NOK 250/hour (approximately US$40) for support agreements (other pricing arrangements available).

• Starting at NOK 250/hour (approximately US$40) for projects or one time jobs. Ranging upwards to NOK 850/hour (approximately US$130).

Last modified:

March 24, 1997.

33.3 UNISON AS

     Holmenveien 20
     Postboks 39, Vindern
     N-0319 Oslo
     Norway

Phone:

+47 22143320 Fax:

+47 22143345 EMail:

info@unison.no URL:

http://www.unison.no Contact:

Per Weisteen <perw@unison.no> Type of support:

Phone, EMail, remote network administration, inhouse and onsite. Special expertise:

10 years of Unix and Internet services in Europe. Training, programming services, porting, installation Internet and intranet infrastructure planning and connectivity, firewall configuration, DNS configuration, sendmail configuration, Usenet News server setup, web server configuration, HTML programming, SQL programming.

Sample prices:

• Starting at NOK 500/hour (approximately US$75) for support agreements

• Starting at NOK 650/hour (approximately US$110) for projects or one time jobs

Last modified:

February 14, 1997.

34 Pakistan

34.1 Vision Systems

     179-C-1 Model Town
     Lahore-54700
     Pakistan

Phone:

+92 42 583 6152, +92 42 586 9353 Fax:

+92 42 5834920 EMail:

abnaseem@paknet1.ptc.pk URL:

http://www.freenet.edmonton.ab.ca/˜ abidn Contact:

Abid Nasim Type of support:

EMail and onsite (Lahore only). Special expertise:

Training, configuration and BBS. Sample prices:

Upon request. Last modified:

March 24, 1997.

35 Poland

35.1 Polskie Sieci Teleinformatyczne S.A.

          ul. Zaolzianska 11
          41-800 Zabrze
          Poland

Phone:

+48 32 176 14 30 Fax:

+48 32 176 01 75 EMail:

pst@netkat.com.pl URL:

http://www.netkat.com.pl Contact:

Andrzej Madloch Type of support:

Phone, EMail, remote network administration, direct help in every main city in Poland. Special expertise:

Training, web/CGI, networking, Internet connectivity planning, design, instalation, configuration and security, application software development.

Sample prices:

US$50-$100/hour dependent on project level. Last modified:

March 15, 1997.

35.2 Publiczny Dostep do Internetu sp. z o.o.

          Matejki 10/1
          87-100 Toru\~{n}
          Poland

Phone:

+48 56 21146 EMail:

rzm@torun.pdi.net URL:

http://www.pdi.net Contact:

Rafal Maszkowski Type of support:

Phone, EMail, remote network administration and direct help in Toruñ area. Special expertise:

Installation, administration, security, connecting to the Internet, training in administration and usage. Sample prices:

Upon request. Last modified:

March 26, 1997.

36 Singapore

36.1 Netstream Pte Ltd

          149, Geylang Road, \#02-03A
          Singapore 389233

Phone:

+65 8429723 Fax:

+65 8429725 EMail:

chlai@netstream.com.sg URL:

http://www.netstream.com.sg Contact:

Lai Chern Horng Type of support:

Phone, EMail, remote network administration, inhouse and onsite. Special expertise:

Intranet and Internet consultants, software development, web services, Internet security consultants. Sample prices:

Upon Request. Last modified:

March 6, 1997.

37 South Africa

37.1 Obsidian Systems

          P.O. Box 4938
          Cresta
          2118
          South Africa

Phone:

+27 11 792 6500 Fax:

+27 11 792 6522 EMail:

info@obsidian.co.za URL:

http://www.obsidian.co.za Contact:

Anton de Wet Type of support:

Onsite. Special expertise:

Installations, training. Mail, web and file server, databases and firewalls. Sample prices:

R150/hour plus VAT, subject to change. Last modified:

March 22, 1997.

37.2 Symphony Research (Pty) Ltd

          P.O. Box 95462
          Grant Park
          2051
          South Africa

Phone:

+27 82 900 8632 Fax:

+27 11 325 0600 EMail:

evan@sr.co.za URL:

http://sr.co.za Contact:

Evan Summers Type of support:

Phone, EMail, remote network administration and onsite. Special expertise:

Internet connectivity with firewalling, web programming, system and network administration. General LAN serving using Linux servers. Four Linux/Internetworking experts, and all C, C++, Java and Perl programmers.

Sample prices:

R200/hour for call outs not covered by maintenance contract. R4000/month per Linux server, for full maintenance and network administration.

Last modified:

March 17, 1997.

38 Spain

38.1 DISAM-SL

     C/ Jos\'{e} Guti\'{e}rrez Abascal 2
     E-28 006 Madrid
     Spain

Phone:

+34 1 336 30 61 Fax:

+34 1 564 29 61 EMail:

marx@disam.upm.es URL:

http://www.disam.upm.es/sl Contact:

Ben Marx, Manuel Escriche Type of support:

Phone, EMail, remote network administration, inhouse and onsite. Special expertise:

Installation, programming services, system administration, Internet and Intranet connectivity, troubleshooting, and training.

Sample prices:

PTS 10.000/hour. Last modified:

March 19, 1997.

39 Sweden

39.1 Guide Datakonsult

          Gustavslundsv. 12
          S-167 51 Bromma
          Sweden

Phone:

+46 70 745 52 39 Fax:

+46 634 46 00 EMail:

mge@guide.se URL:

http://wwwsto.guide.se Contact:

Mikael Geijer Type of support:

EMail, inhouse. Special expertise:

Internet services, Internet payment methods, Java applications and development. Sample prices:

Upon request. Last modified:

February 19, 1997.

39.2 Signum Support AB

     Teknikringen 8
     S-583 30 Link\"{o}ping
     Sweden

Phone:

+46 13 21 46 00 Fax:

+46 13 21 47 00 EMail:

info@signum.se URL:

http://www.signum.se Contact:

Magnus Redin <redin@signum.se> Type of support:

Phone, EMail, fax and remote network administration. Special expertise:

Development of the Fuego firewall sold as a complete package including support. Consulting for the industrial market and development of customized Linux kernels for embedded systems. Readynet Internet server, Readycom conferensing system and Phttpd support, all three are in house developed free software.

Sample prices:

• US$110/hour for consulting

• Approximately US$4300 for Readynet Internet server installation and support

Last modified:

March 16, 1997.

40 Turkey

40.1 Hakan Kalyoncu

     Perihan S 49/12
     Sisli 80260 Istanbul
     Turkey

Phone:

+90 532 323 4186 EMail:

hk@yore.com.tr URL:

http://www.yore.com.tr/˜ hk Contact:

Hakan Kalyoncu Type of support:

Phone, EMail, inhouse and onsite. Special expertise:

Training, Internet connectivity, Internet server setup, administration. Sample prices:

• Free for individuals

• US$100/hour for corporates

Last modified:

February 19, 1997.

41 United Kingdom

41.1 Cygnet Internet Services Ltd

          Lee Valley Technopark
          Ashley Road
          London, N17 9LN
          UK

Phone:

+44 181 880 4650 Fax:

+44 181 880 4651 EMail:

steve@cygnet.co.uk URL:

http://www.cygnet.co.uk Contact:

Steve Merrony Type of support:

Phone, EMail and inhouse. Special expertise:

All aspects of Linux system administration, news server configuration, web and database systems. Sample prices:

Consultancy from £400/day. Last modified:

March 15, 1997.

41.2 Paul Dunne

          30 Onslow Gardens
          London, N10 3JU
          UK

Phone:

+44 181 374 8194 EMail:

paul@tiny1.demon.co.uk URL:

http://www.tiny1.demon.co.uk Contact:

Paul Dunne Type of support:

EMail and onsite. Special expertise:

Linux installation and configuration, Internet connectivity, system administration, programming with shell and Perl scripts.

Sample prices:

Basic rate £30/hour. Last modified:

March 15, 1997.

41.3 Laboratory for Micro Enterprise

          Oakley Court
          Benson
          Wallingford, OX10 6QH
          UK

Phone:

+44 1491 832932 Fax:

+44 1491 832932 EMail:

lme@storrs.demon.co.uk URL:

http://www.i-way.co.uk/˜ storrs Contact:

John Storrs Type of support:

As required. Special expertise:

Our focus is to support very small enterprises (most of us) with affordable systems and tools. See our web site for details.

Sample prices:

Upon request. Last modified:

March 15, 1997.

41.4 Marketgrid Limited

          Britannia Road
          Poole
          Dorset, BH14 8AZ
          UK

Phone:

+44 1202 731056 EMail:

philip.smith@marketgrid.com URL:

http://www.marketgrid.com Contact:

Philip Smith Type of support:

Phone, EMail, remote network administration, inhouse and onsite. Special expertise:

Programming: C, C++, Java, SQL; Linux installation and administration. Network setup and administration, Internet development: HTML, CGI, ActiveX, Java. Databases: SQL server, Sybase, C-Tree.

Sample prices:

Consultancy from £400/day. Last modified:

April 9, 1997.

41.5 Nene Information Sytems

     32 Glasgow Street
     Northampton, NN5 5BL
     UK

Phone:

+44 1604 464561 Fax:

+44 1604 464561 EMail:

sales@nisys.demon.co.uk URL:

http://www.nisys.demon.co.uk Contact:

Timothy Peters Type of support:

EMail and inhouse.

Special expertise:

Software development (Java, C, C++, Perl), web page design, network inatallations, Internet and network connectivity solutions, pre-installed Linux servers (Intel and Alpha), Cyclades multiport serial cards, Linux distribution CD-ROMs, Books and other software.

Sample prices:

£25/hour for software development. Last modified:

March 19, 1997.

42 USA - General

42.1 InfoMagic, Inc

     11950 N. Highway 89
     Flagstaff, AZ 86004
     U.S.A.

Phone:

1 900 786 5555 Fax:

1 520 526 9573 EMail:

support@infomagic.com URL:

http://www.infomagic.com/support/linux Contact:

Henry M. Pierce Type of support:

Individuals who purchase InfoMagic Linux Developer’s Resource get free EMail and fax back support. Support via phone for our Linux Developer’s Resource or any Linux vendor’s Linux CD is US$2.00/minute in the U.S.A. Everyone is welcome to look at our support web pages. International customers needing phone support should contact us via EMail first.

Special expertise:

Everything related to InfoMagic Linux Developer’s Resource. Sample prices:

Fixed service contracts, including remote maintance are available starting at US$50/engineering hour. Please contact us by EMail for additional information.

Last modified:

February 19, 1997.

42.2 NDA

          400 West Cummings Park
          Woburn, MA 01801
          U.S.A.
          650 Castro Street, Suite 120-209
          Mountain View, CA 94041-2012
          U.S.A.

Phone:

+1 617 937 3338 Fax:

+1 617 937 3775 EMail:

info@nda.com URL:

http://www.nda.com Contact:

Tom Bills, Mary Ede Type of support:

Phone, EMail, remote network administration and onsite. Special expertise:

Unix. Sample prices:

US$50-$150 based on task. Last modified:

March 15, 1997.

42.3 PSA AnswerDesk – Pencom Systems Administration

     9050 Capital of Texas Hwy North, Suite 310
     Austin, TX 78759
     U.S.A.

Phone:

+1 512 343 1111 or contact your nearest Pencom office. Fax:

+1 512 346 6444 EMail:

answer-info@pencom.com URL:

http://www.pencom.com/psa/answer.html Contact:

Mike Carpenter Type of support:

Phone. Special expertise:

AnswerDesk is an Open Systems phone support desk specifically designed to provide a backstop with difficult, senior administration issues. We do not provide end user support. We support administrators working with all Unix platforms, including Linux. Our client list includes a wide variety of Fortune 500 companies utilizing open systems, many with mission critical support requirements. Our technical database is supported by 200 of the best systems administrators in the business.

Sample prices:

AnswerDesk charges US$100/hour, with a one hour minimum. If we don’t solve your problem, you are not charged. Please request a brochure for more in-depth info.

Last modified:

March 15, 1997.

43 USA - Arizona

43.1 M. Cooper

     P.O. Box 237
     St. David, AZ 85630
     U.S.A.

Phone:

+1 520 720 9431 EMail:

thegrendel@theriver.com URL:

http://personal.riverusers.com/˜ thegrendel Contact:

Leo M. Cooper Type of support:

Phone and EMail. Special expertise:

We offer setup, programming and training. Sample prices:

US$20-$50/hour, negotiable. Last modified:

March 15, 1997.

43.2 Mechanix Computer Consulting

          1515 S. Extension
          Mesa AZ 85210
          U.S.A.

Phone:

+1 603 668 6325 Fax:

+1 603 668 6325 EMail:

shindle@toocool.com, shindle@goodnet.com Contact:

Stephen Hindle Type of support:

Phone, EMail, remote network administration and onsite.

Special expertise:

15 years programming, experience with Linux kernels from 0.99 to 2.1.x. Installation, programming services, Perl and connectivity.

Sample prices:

• US$40/hour for programmer analyst

• US$35/hour for programmer

• Hourly rates vary with length and nature of contract

• Fixed price contracts are also available

Last modified:

March 15, 1997.

43.3 M&S Group, Inc.

          1038 E Brentrup Dr.
          Tempe, AZ 85283
          U.S.A.

Phone:

602-491-7599 Fax:

602-491-7599 EMail:

nikitin@asu.edu URL:

http://www.m-and-s.com Contact:

Sergey Nikitin Type of support:

Phone, EMail, inhouse and onsite. Special expertise:

Training, workstation design and manufacturing. Sample prices:

Upon request. Last modified:

March 19, 1997.

44 USA - California

44.1 Evolve Computer Solutions

          1556 Rambla Brisa Drive
          San Marcos, CA 92069
          U.S.A.

Phone:

+1 619 746 2757 Fax:

+1 619 751 0612 EMail:

info@evolveinc.net URL:

http://www.evolveinc.net Contact:

Sean P. Kane Type of support:

Phone, EMail and inhouse. Special expertise:

System design (Complete hardware and software solutions), installation (Distributions, X11, applications), etc.

Sample prices:

US$40-$75/hour. Last modified:

March 18, 1997.

44.2 Inter@ctivate Consulting Group

          11172 Morning Creek Drive, South
          SanDiego, CA 92128
          U.S.A.

Phone:

+1 619 748 7684 Fax:

+1 619 748 7684 EMail:

interact@interactivate.com URL:

http://www.interactivate.com/ Contact:

Sean Dreilinger Type of support:

Phone, EMail, remote network administration and onsite. Special expertise:

Research, training, and marketing of networked information services, with an eye towards PD and GPL solutions, i.e. Linux!

Sample prices:

Free consultation. Per project and hourly rates negotiable. Last modified:

February 20, 1997.

44.3 Linux-Consulting

          RING, Inc
          P.O. Box 70958
          Sunnuyvale, CA 94086
          U.S.A.

Phone:

+1 408 245 8400 Fax:

+1 408 523 9097 EMail:

alvin@planet.fef.com, alvin@Linux-Consulting.com URL:

http://planet.fef.com, http://www.linux-consulting.com Contact:

Alvin Oga Type of support:

EMail, remote network administration and HOWTO-miniSearchEngine.

Special expertise:

We install, setup, test and maintain your Linux machines, provide custom Internet access and connectivity and custom Internet servers.

Sample prices:

Upon request. Last modified:

March 16, 1997.

44.4 RCT Design

     Santa Barbara, CA 93117
     U.S.A.

Phone:

+1 805 683 9717 Fax:

+1 805 683 0717 EMail:

bob@zooid.com URL:

http://www.zooid.com Contact:

Bob Tellefson Type of support:

Phone, EMail and remote network administration. Special expertise:

Turnkey Linux systems with support, Internet and Intranet servers, remote access, gateways, firewalls, web site design, database access, and CGI programming.

Sample prices:

US$45/hour. Last modified:

March 19, 1997.

44.5 RG Consulting

          1751 East Roseville Pkwy. \#1828
          Roseville, CA 95661
          U.S.A.

Phone:

+1 916 786 7945 Fax:

+1 916 786 5311 EMail:

rfg@monkeys.com Contact:

Ronald F. Guilmette Type of support:

Inhouse and onsite. Special expertise:

Maintenance, and enhancement of GNU software development tools. Sample prices:

Upon request. Last modified:

February 19, 1997.

44.6 Bill Rousseau

          842 Lucille Street
          Livermore, CA 94550-3519
          U.S.A.

Phone:

+1 510 455 8008 Fax:

+1 510 455 8008 EMail:

rousseau@aimnet.com URL:

Look me up in http://www.ieee-sv-consult.org which I maintain.

Contact:

Bill Rousseau Type of support:

Phone, EMail and onsite. Special expertise:

Software development is offsite as much as possible. Training is available if there is sufficient interest. Sample prices:

General computing consulting (Unix, Linux, network configuring, software development) US$100/hour. Support for individuals attempting to install or configure Linux for their own use or for work that promises to benefit the Linux community may be offered at much lower prices or for free in special cases.

Last modified:

March 16, 1997.

44.7 Softcraft Impresa

     2923 Cohansey Drive
     San Jose, CA 95132
     U.S.A.

Phone:

+1 408 251 9820 EMail:

mauro@softcraft.com URL:

http://www.softcraft.com Contact:

Mauro DePalma Type of support:

Linux Generations (see below) installs are supported, within reasonable expectations, for one year, at no additional cost.

Special expertise:

We can fine tune homogeneous and heterogeneous networks with Linux/Unix nodes and others TCP/IP capable. We also have an evolving secure operating environment based on Linux which is complete but does not (yet) come with installation scripts/programs and is currently only available for Intel. We call it: Linux Generations. Before any other distributions, we know of, shadow passwords and groups work seamlessly with NYS/DNS for secure nodes, gateways and networks. Additionally, with the use of Samba (SMB protocol) shares to and from Windows 95, Windows NT and Windows 3.x with TCP/IP are easy to setup. Talk to us about Java and CGI, coupled with the Apache HTTP server this is the way to have server-side Java: the power of CGI expressed with Java applications.

Sample prices:

US$90/hour. Depends on the kind of work and duration. Last modified:

March 15, 1997.

44.8 Earl A. Stutes, Inc.

     P.O. Box 35567
     Monte Sereno, CA 95030-0567
     U.S.A.

Phone:

+1 408 448 1089 Fax:

+1 408 448 7864 EMail:

estutes@eas.san-jose.ca.us URL:

http://www.eas.san-jose.ca.us Contact:

Earl Stutes Type of support:

Phone, EMail and inhouse. Special expertise:

Web engineering, Internet security, Perl and C++ programming. Sample prices:

Upon Request. Last modified:

March 15, 1997.

44.9 Vennerable Consultants

          7216 Lincoln Avenue, Suite 101
          El Cerrito, CA 94530-3240
          U.S.A.

Phone:

+1 510 525 2989 Fax:

+1 510 525 8072 EMail:

jason@vennerable.com URL:

http://www.vennerable.com Contact:

Jason Venner Type of support:

Phone, EMail, remote network administration and inhouse. Special expertise:

Linux, FreeBSD, X11 (servers and applications), C, C++, Perl, tcl/tk, web support services, secure web servers and distributed systems.

Sample prices:

Upon reqest. Last modified:

February 19, 1997.

44.10 Gregg Weber

          1076 Carol Lane \#3
          Lafayette, CA 94549
          U.S.A.

Phone:

+1 510 283 6264 EMail:

gregg@netcom.com Contact:

Gregg Weber Type of support:

Onsite in SF bay area, or can travel for big jobs. Special expertise:

Installation, C programming services, UUCP for netware networks, troubleshooting. Sample prices:

US$55/hour. Last modified:

February 19, 1997.

44.11 Yggdrasil Computing, Inc.

     4880 Stevens Creek Boulevard, Suite 205
     San Jose, CA 95129-1034
     U.S.A

Phone:

+1 800 261 6630 Fax:

+1 408 261 6631 EMail:

tech-support@yggdrasil.com URL:

http://www.yggdrasil.com Contact:

Bill Selmeier Type of support:

Phone, EMail and remote network administration. Special expertise:

We are also able to use our T1 Internet connection to solve your problem online (if your machine is also on the Internet).

Sample prices:

US$25 for a 15 minute support call guaranteed to solve your problem. Other arrangements are available.

Last modified:

February 19, 1997.

45 USA - Connecticut

45.1 BRT Technical Services Corporation

          50 Newtown Road
          Danbury, CT 06810
          U.S.A.

Phone:

+1 203 748 5100 Fax:

+1 203 792 0280 EMail:

info@brttech.com URL:

http://www.brttech.com Contact:

Chad Robinson Type of support:

Phone, EMail and inhouse. Special expertise:

Training. Sample prices:

Upon reqest. Last modified:

February 20, 1997.

45.2 Creative Systems

          Clearview Knoll
          Middlebury, CT 06762
          U.S.A.

Phone:

+1 203 598 3690 Fax:

+1 203 377 7931 EMail:

creative@stratgrp.com URL:

http://www.creative.stratgrp.com Contact:

Silas Moeckel Type of support:

Phone, EMail, remote network administration, inhouse and onsite. Special expertise:

Installation, programming services, system administration, Internet connectivity, troubleshooting and training.

Sample prices:

• US$100/hour for troubleshooting (Four hour minimum) US$50/hour for programming and porting

Last modified:

February 25, 1997.

45.3 Darcom Systems Ltd

          Suite 6, 11 Wall St
          Milford, CT 06460
          U.S.A.

Phone:

+1 203 878 6640 EMail:

info@darcom.dyndns.com URL:

http://darcom.dyndns.com Contact:

Peter D’Arco Type of support:

Phone, EMail, remote network administration, inhouse and onsite. Special expertise:

Installation, system administration, Internet and Intranet connectivity, firewall installation, troubleshooting, training, complex mail systems, DNS, NIS, network infrastructure.

Sample prices:

US$50-$150/hour. Last modified:

March 16, 1997.

45.4 Kracked Rock Komputing, LLC

     13 Fleetwood Drive
     Danbury, CT 06810-7009
     U.S.A.

Phone:

+1 203 778 6269 Fax:

+1 203 778 6292 EMail:

ken@eci.com Contact:

Kenneth E. Nawyn Type of support:

Phone, EMail and inhouse. Special expertise:

systems administration, system configuration and technical support for Linux and FreeBSD. Over ten years of Unix systems adminstration experience along with Perl and C coding experience. For more detailed description of our experience and services, please contact us.

Sample prices:

Upon request. Last modified:

March 27, 1997.

46 USA - Colorado

46.1 Colorado Computer Consultants

  615 Quartz Way
  Broomfield, CO 80020-1727
  U.S.A.

Phone:

+1 303 404 9589 Fax:

+1 303 404 9588 EMail:

co3-info@co3.com URL:

http://www.co3.com Contact:

Robert Kaczanowsk Type of support:

Phone, EMail and onsite. Special expertise:

Installation. Sample prices:

• US$35/hour for individuals (appointment required)

• US$55/hour for individuals (immediate dispatch)

• US$45/hour for small business (appointment required)

• US$65/hour for small business (immediate dispatch)

• US$55/hour for corporates (appointment required)

• US$100/hour for corporates (immediate dispatch)

• Contract prices available.

May be subject to additional charges:

• Mileage: US$0.26 per mile over 5 miles

• Parking as necessary

• Permits and fees as necessary

• Additional supplies as necessary

Last modified:

March 18, 1997.

46.2 Daylight Software

          1062 Lexington Lane
          Estes Park, CO 80517
          U.S.A.

Phone:

+1 970 685 6058 EMail:

daylight@frii.net URL:

http://www.frii.net/˜ daylight Contact:

Chris Howard Type of support:

Phone and remote network administration. Special expertise:

Administration and programming. Sample prices:

US$65/hour. Last modified:

April 2, 1997.

46.3 Granite Computing Solutions

          P.O. Box 270103
          Fort Collins, CO 80527-0103
          U.S.A.

Phone:

+1 970 225 2370 EMail:

granite@SoftHome.Net URL:

http://www.SoftHome.Net/granite Contact:

Brian Grossman <brian@SoftHome.Net> Type of support:

Phone, EMail, remote network administration and onsite. Special expertise:

Training developed on request for Internet and Unix topics. Sample prices:

US$45/hour for technical support. Last modified:

February 19, 1997.

46.4 SNLutz Associates

          3817 Cottage Drive
          Colorado Springs, CO 80920
          U.S.A.

Phone:

+1 719 590 4210 EMail:

SNLutz@aol.com Contact:

Scott Lutz Type of support:

Phone and EMail. Special expertise:

Programming services. Sample prices:

US$50/hour. Last modified:

April 6, 1997.

46.5 Worldwide Solutions, Inc.

          4450 Arapahoe Ave., Suite 100
          Boulder, CO 80303
          U.S.A.

Phone:

+1 303 581 0800 Fax:

+1 303 530 0191 EMail:

info@wwsi.com URL:

http://www.wwsi.com Contact:

Steve Hultquist, Bob Davenport Type of support:

Phone, EMail, remote network administration and onsite. Special expertise:

Training, installations, configuration, specification, architecture, engineering, integration. Sample prices:

Variable. Monthly support agreements or project based statement of work pricing preferred. Standard hourly rates available for those with specific requirements.

Last modified:

March 26, 1997.

47 USA - Florida

47.1 LOD Communications

          1621 Creighton Road
          Pensacola, FL 32504
          U.S.A.

Phone:

+1 904 494 0955 EMail:

support@lod.com URL:

http://www.lod.com/cns.html Contact:

Todd Lawrence <lawrence@lod.com> Type of support:

Phone, remote network administration and onsite. Special expertise:

We provide training in all forms of Unix system administration and networking. We can generally provide our own transportation anywhere in the continental USA for onsite consultation.

Sample prices:

• US$30 for phone support

• US$50/hour with fixed/flat rate estimates for remote (telnet/rlogin) consulting

• US$75/hour with fixed/flat rate estimates for onsite Unix consultation

Last modified:

February 19, 1997.

47.2 Progressive Computer Concepts, Inc.

          1371 Cassat Avenue
          Jacksonville, FL 32205
          U.S.A.

Phone:

+1 800 580 2640 Fax:

+1 904 389 6584 EMail:

info@progressive-comp.com URL:

http://www.progressive-comp.com Contact:

Lester Hightower Type of support:

Phone, EMail, remote network administration and onsite.

Special expertise:

Systems programming, applications programming, system installation and administration (both remote and onsite), Internet and Intranet connectivity, LAN and WAN, firewall installation, security, troubleshooting and training, Unix accounting software, Samba SMB networking support, custom LAN/WAN/Intranet business systems development, WWW/CGI development (e.g. database gateways, catalogs).

Sample prices:

• US$52/hour offsite, $60 onsite for installation and administration

• US$60/hour for programming

• US$52/hour offsite, $60 onsite for LAN/WAN consulting

• US$60/hour for Internet and Intranet

• US$45/hour for EMail support

Last modified:

February 20, 1997.

48 USA - Georgia

48.1 Fly-By-Day Consulting, Inc.

     4642 Bentley Place
     Duluth, GA 30136
     U.S.A.

Phone:

+1 770 662 8321 Fax:

Upon request. EMail:

bob@cavu.com URL:

http://www.mindspring.com/˜ cavu Contact:

Bob Toxen Type of support:

Phone, EMail, fax, remote network administration, inhouse and onsite.

Special expertise:

Porting, developing client-server applications, system administration, networking (Internet and Intranet), web setup, file system repair, Linux setup. 22 years of UNIX and C experience, reasonable rates.

Sample prices:

Upon request. Last modified:

March 18, 1997.

48.2 Triangle Software Corporation

     3216 Caldwell Road
     Atlanta, GA 30319
     U.S.A.

Phone:

+1 800 625 7787, +1 770 455 9089 Fax:

Upon request. EMail:

triangle@america.net Contact:

Ben Cooper Type of support:

Inhouse. Special expertise:

Inhouse ANSI C++ training with labs using G++ (very well reviewed by students). Sample prices:

US$5000 plus travel expenses for a five-day C++ class with up to 15 students, taught at your site. Last modified:

March 16, 1997.

49 USA - Illinois

49.1 Rahim Azizarab

          1925 S 3rd Ave
          Maywood, IL 60153
          U.S.A.

Phone:

+1 708 344 6994 EMail:

rahim@megsinet.net or ray@tezcat.com Contact:

Rahim Azizarab Type of support:

Phone, EMail and inhouse. Special expertise:

Installation, programming services, system administration, Internet and Intranet connectivity, firewall installation, troubleshooting and training.

Sample prices:

Free for individuals. Last modified:

February 22, 1997.

49.2 onShore, Inc.

          1407 W. Chicago
          Chicago, IL 60622
          U.S.A.

Phone:

+1 773 667 4673 Fax:

+1 773 667 4263 EMail:

sales@onshore.com URL:

http://www.onshore.com Contact:

Stelios Valavanis Type of support:

Phone, EMail and inhouse. Special expertise:

Inhouse installation and support of Linux for custom solutions such as routing, firewall, EMail, web serving, etc. Custom solutions for job tracking, print spooling/processing, Intranet, etc.

Sample prices:

• US$110/hour for support

• US$125/hour for software development

Last modified:

March 18, 1997.

50 USA - Indiana

50.1 Citadel Software, Inc.

          241 East Eleventh Street
          P.O. Box 407
          Brookville, IN 47012-0407
          U.S.A.

Phone:

+1 765 647 4720 Fax:

+1 765 647 6059 EMail:

lfrost@cnz.com URL:

http://www.cnz.com Contact:

Lyle Frost Type of support:

Remote network administration and onsite.

Special expertise:

Citadel provides a complete line of professional computer services, including network and Intranet design, installation, maintenance, and administration, as well as software development, training, and technical writing.

Sample prices:

Upon request. Last modified:

March 24, 1997.

50.2 LANscape

     1100 W Lloyd Expressway
     Evansville, IN 47708
     U.S.A.

Phone:

+1 800 896 6171, +1 812 423 1231 Fax:

+1 812 426 6138 EMail:

info@lanscape.net URL:

http://www.lanscape.net Contact:

Steven Pritchard Type of support:

Limited free phone and EMail support available for Linux users as time permits. Ongoing phone and/or EMail support available with a service contract.

Special expertise:

TCP/IP network setup and maintenance, system administration, programming, and training available on-site or at/from our office. We also offer web site hosting, POP mail accounts, EMail forwarding, primary and secondary DNS, etc. See our web site for details.

Sample prices:

Base rate is US$100/hour for most services. Contract pricing is available at a reduced rate. Last modified:

April 1, 1997.

51 USA - Maine

51.1 Midcoast Internet Solutions

          P.O. Box 669
          Rockland, ME 04841
          U.S.A.

Phone:

+1 207 594 8277 Fax:

+1 207 596 7248 EMail:

webmaster@midcoast.com URL:

http://www.midcoast.com/mis Contact:

Jason Philbrook Type of support:

Phone, EMail, remote network administration and onsite. Special expertise:

Connectivity, troubleshooting, system setups. Sample prices:

US$50/hour minimum. Last modified:

March 12, 1997.

52 USA - Maryland

52.1 I-Link, Inc.

          136 Forests Edge Place,
          Laurel, MD 20724
          U.S.A.

Phone:

+1 301 604 8411 Fax:

+1 301 604 8411 EMail:

fbennett@i-linkcom.com URL:

http://i-linkcom.com Contact:

Fred Bennett Type of support:

Phone, EMail and inhouse. Special expertise:

We support software from Red Hat Software, InfoMagic, Walnut Creek, and Caldera. We are an Authorized Caldera Channel partner.

Sample prices:

US$50-$100/hour. Last modified:

March 16, 1997.

52.2 Matthew Kaylor

          13639 Barnhart Rd.
          Clear Spring, MD 21722
          U.S.A.

Phone:

+1 301 733 2922 EMail:

maxit@nfis.com URL:

http://www.nfis.com/˜ maxit name = ”http://maxit.nfis.com”¿ Contact:

Matthew Kaylor Type of support:

Phone, EMail, IRC and world wide travel for onsite support.

Special expertise:

Linux system setup (remote or inhouse) to commercial networked Internet servers (complete). Personal or group Linux training courses.

Sample prices:

• US$5000/month, US$300/day, US$40/hour

• Negotiated pricing based on needs for individuals

Last modified:

February 20, 1997.

52.3 Bradley M. Kuhn

          Baltimore, MD
          U.S.A.

Phone:

+1 410 662 5846 EMail:

bkuhn@acm.org URL:

http://www.smart.net/˜ bkuhn/resume.shtml Contact:

Bradley M. Kuhn Type of support:

Phone, EMail and remote network administration. Special expertise:

Installation, integration, Perl developement, system administration. Sample prices:

US$35-$45/hour depending on work. Substantially less for non-profit organizations. Last modified:

March 5, 1997.

52.4 Vertex Development

     11 Crestview Court
     Potomac, MD 20854
     U.S.A.

Phone:

+1 301 340 2162 Fax:

+1 301 309 6968 EMail:

jeff@vertexdev.com URL:

http://www.vertexdev.com Contact:

Jeff Barr Type of support:

Phone, EMail and inhouse. Special expertise:

Porting Unix and non-Unix software to Linux. User Interface design, development, and review. Applications and systems software development.

Sample prices:

Porting and development projects are typically handled on a firm-fixed price basis. Other services are supplied at US$120-$150/hour, depending on duration and travel requirements.

Last modified:

March 17, 1997.

53 USA - Massachusetts

53.1 Working Version

     31 Shea Road
     Cambridge, MA 02140
     U.S.A.

Phone:

Scheduled phone calls by appointment; please use EMail generally.

Fax:

Scheduled fax reception by appointment; please EMail PostScript generally.

EMail:

bb@wv.com

URL:

http://www.wv.com

Contact:

Brian Bartholomew

Type of support:

Phone, EMail, fax, remote network administration, inhouse and onsite.

Special expertise:

We generally do contract program development, Unix toolsmithing, Unix piecework, and version-control infrastructure consulting, but we’re very flexible and highly negotiable. Here are some keywords to describe our expertise base: configuration management, version control, ClearCase, telco, tariffs, Internet service provider, move planning and execution, ISP cutover, vendor management, systems architect, systems analysis and planning, systems design, feasibility and cost benefit studies, business process reengineering, enterprise integration, technology infrastructure requirements determination, project planning and scheduling, prototyping and evaluation, requirements determination, toolsmithing, network design and implementation, network monitoring, network mapping, network visualization, facilities planning, machine room layout, data center design, automated regression testing, build systems, financial, trading floor, distributed systems, database design, Unix system administration.

Sample prices:

US$65-$250/hour. Please EMail requirements for a quote on fixed-price development projects.

Last modified:

March 17, 1997.

54 USA - Michigan

54.1 Jim Willette

     18190 - 15 Mile Rd
     Big Rapids, MI 49307
     U.S.A.

Phone:

+1 616 592 3119 EMail:

jimw@ctec.net URL:

http://www.netonecom.net/˜ jimw, http://www.ctec.net Contact:

Jim Willette Type of support:

Phone, EMail, remote network administration and inhouse. Special expertise:

Specialties include Internet firewalls, proxy servers, web servers, domain name servers (DNS), Linux terminal servers, Internet connectivity networking, NFS and more.

Sample prices:

• US$45/hour for phone support

• US$55/hour for onsite plus expenses and travel time (negotiable)

Last modified:

March 27, 1997.

55 USA - Missouri

55.1 Reboot Inc.

          579 Rue St. Ferdinand
          Florissant, MO 63031
          U.S.A.

Phone:

+1 314 830 0427 Fax:

+1 314 830 0427 EMail:

reboot@cyberusa.com URL:

http://www.cyberusa.com/˜ reboot Contact:

Zubin Chandran Type of support:

Phone, EMail, and onsite to the 314 and 618 area codes. Special expertise:

Training onsite as well as hardware upgrades and sales. Sample prices:

Call for special pricing and promotional deals.

• US$60 for up to the first hour and 15 minutes, then US$15 every 15 minutes thereafter for onsite support.

• Phone support is 60 cents per minute.

Last modified:

March 17, 1997.

55.2 Damon Richardson

     St. Louis, MO
     U.S.A.

EMail:

nomadstl@concentric.net URL:

http://www.concentric.net/˜ nomadstl/linux.htm Contact:

Damon Richardson Type of support:

EMail. Special expertise:

Internet web server setup, EMail POP3 server setup, dial-up server setup, file and print sharing setup with Windows and Novel clients. Installation, programming services, system administration, Internet and Intranet connectivity, and training.

Sample prices:

Upon request. Last modified:

February 17, 1997.

56 USA - Nebraska

56.1 Probe Technology Inc

     10159 J St
     Omaha, NE 68164
     U.S.A.

Phone:

+1 402 593 9800 Fax:

+1 402 593 8748 EMail:

info@probe.net URL:

http://www.probe.net Contact:

Dale Botkin, Tim Russell Type of support:

Phone and EMail. Special expertise:

Training, configuration. Can configure Internet routers, firewalls, firewalls, mail servers, web servers (Apache), etc. Can also provide hardware platform of choice at competitive prices, as well as provide a near-turnkey solution for an ISP setup. Training available.

Sample prices:

US$100/hour consulting, average setup of new system with services is 5-7 hours. Last modified:

March 17, 1997.

56.2 tummy.com, ltd.

     3506 Stratton Drive
     Fort Collins, CO 80525-2722
     U.S.A.

Phone:

+1 970 223 8215 Fax:

+1 408 490 2728 EMail:

info@tummy.com URL:

http://www.tummy.com Contact:

Sean Reifschneider <jafo@tummy.com> Type of support:

Phone, EMail, inhouse. Special expertise:

Custom programming, firewall setup, general Unix/Linux consulting. Sample prices:

Depending on level of expertise, usually in the US$60-$125 range. Last modified:

February 20, 1997.

57 USA - Nevada

57.1 Steve Hedlund

          Las Vegas, NV
          U.S.A.

Phone:

+1 702 254 6749 EMail:

ssh@nevadacom.com URL:

http://nevadacom.com Contact:

Steve Hedlund Type of support:

Phone, EMail and inhouse.

Special expertise:

Internet and Intranet connectivity, installation, programming services, system administration, firewall installation, troubleshooting, virtual addressing, IP Masquerading (for connecting networks to the Internet).

Sample prices:

US$25-$40/hour. Last modified:

March 19, 1997.

58 USA - New Jersey

58.1 Etc Services

     31 N. Vivyen St.,
     Bergenfield, NJ 07621-1529
     U.S.A.

Phone:

+1 201 385 7113 EMail:

swebster@carroll.com Contact:

Scott D. Webster Type of support:

Inhouse. Special expertise:

System setup and administration, TCP/IP network configuration, web server setup (not content creation), Linux LAN servers, DNS configuration, experience with Linux on x86, Linux on Digital Alpha (AXP).

Sample prices:

Upon request. Last modified:

March 17, 1997.

58.2 ZEI Software

          2713 State Hwy 23 South
          Newfoundland, NJ 07435
          U.S.A.

Phone:

+1 201 208 8800 Fax:

+1 201 208 1888 EMail:

art@nji.com Contact:

Arthur Zysk Type of support:

Phone, EMail, inhouse and onsite. Special expertise:

Senior project managers with 25 years of experience. C, C++, Sybase, kernel drivers, Internet Connectivity. Mission critical project management services.

Sample prices:

Upon request. Last modified:

February 19, 1997.

59 USA - New York

59.1 Crynwr Software

          521 Pleasant Valley Rd.
          Potsdam, NY 13676
          U.S.A.

Phone:

+1 315 268 1925 Fax:

+1 315 268 9201 EMail:

info@crynwr.com URL:

http://www.crynwr.com Contact:

Russell Nelson Type of support:

Familiar with most topics. Special expertise:

Qualified to do any kind of support; specializing in networking drivers. Sample prices:

US$100/hour; fixed-price quotes for your task. Last modified:

March 15, 1997.

60 USA - Ohio

60.1 DESIGN SCIENCE LABS

          P.O. Box 542
          Berea, OH 44017-0542
          U.S.A.

Phone:

+1 216 243 1346 Fax:

+1 216 524 0979 EMail:

stutz@dsl.org URL:

http://www.dsl.org Contact:

Michael Stutz Type of support:

Phone, remote network administration and onsite. Special expertise:

We are exclusive to Linux and GNU and provide comprehensive hypermedia/Internet solutions including web design, analog/digital logo/illustration, technical writing (etc. TeX, texinfo, SGML) and training.

Sample prices:

Billing by project or hourly. US$50/hour for most work. Last modified:

March 16, 1997.

60.2 Internet Robotics

     131 Willow Lane, Floor 2
     Oxford, OH 45056
     U.S.A.

Phone:

+1 513 523 7621 Fax:

+1 513 523 7501 EMail:

support@bill.iac.net URL:

http://www.minivend.com Contact:

Mike Heins Type of support:

Phone and EMail, will travel for training courses. Special expertise:

Specializing in online shopping with MiniVend, the GPL shopping cart. Expertise in all aspects of Linux, web administration, shell and Perl scripting and network configuration. Performance tuning.

Sample prices:

• US$75/hour for contract shopping site HTML design

• US$100 for monthly retainer (includes first hour)

• US$85/hour for contract technical support (with retainer)

• US$85/hour for contract shopping software design

• US$100/hour for expert technical support (uncontracted)

Last modified:

February 19, 1997.

60.3 Practical Network Design

     9 Chambers Rd
     Mansfield, OH 44906-1302
     U.S.A.

Phone:

+1 419 529 3841 Fax:

+1 419 529 3625 EMail:

info@practical.net URL:

http://www.practical.net Contact:

Chuck Stickelman Type of support:

Phone, EMail and onsite. Special expertise:

PND specializes in Unix-based solutions, with a strong emphasis on Linux. We provide system and network planning, sales, installation, support and administration. Custom programming and application development is also available, as well as training for system and network administration. Hardware service and support is available to clients in the North-Central Ohio area.

Sample prices:

• US$50/hour for phone, dial-up and Internet support (One hour minimum)

• Onsite support can be contracted on an hourly basis at US$100/hour (Two hour minimum) or by half-day/full-day or per project basis.

Last modified:

March 18, 1997.

61 USA - Oregon

61.1 Kymsoft Consulting

          745 Foothill Dr
          Eugene, OR 97405
          U.S.A.

Phone:

+1 541 465 9008 Fax:

+1 541 465 9052 EMail:

rick@efn.org URL:

http://www.efn.org/˜ rick Contact:

Rick Bronson Type of support:

Phone, EMail and inhouse. Special expertise:

12 years Unix experience. Sample prices:

US$65/hour. Last modified:

February 19, 1997.

61.2 NET-Community

          522 SW 5th Avenue
          Suite 1105
          Portland, OR 97204
          U.S.A

Phone:

+1 800 919 0060, +1 503 274 4423 Fax:

+1 503 953 5829 EMail:

sales@net-community.com URL:

http://www.net-community.com Contact:

Scott Christley Type of support:

Phone and EMail. Special expertise:

NET-Community specializes in support for GNUstep and MediaBook. Our web site has details regarding the support plans.

Sample prices:

• US$50/year for GNUstep Developer (EMail only)

• US$200/year for GNUstep Corporate (EMail only)

• US$2500/year for MediaBook Comprehensive (EMail and phone)

Last modified:

March 18, 1997.

62 USA - Pennsylvania

62.1 Christopher J. Fearnley

          4807 Chester AVE APT 207
          Philadelphia, PA 19143-3441
          U.S.A.

Phone:

+1 215 724 2265 EMail:

cjf@netaxs.com URL:

http://www.netaxs.com/˜ cjf Contact:

Chris Fearnley Type of support:

Phone, ssh-based remote network administration and onsite. Special expertise:

Debian GNU/Linux configurations for TCP/IP networks. Installation and maintenance of various types of Internet servers including web, news, and mail. Business and technical consulting for ISPs.

Sample prices:

US$500/day. US$65/hour. Last modified:

February 19, 1997.

63 USA - Texas

63.1 xprt Computer Consuting, Inc.

          731 Voyager
          Houston, TX 77062
          U.S.A.

Phone:

+1 281 480 8649 Fax:

+1 281 480 7357 EMail:

jody@sccsi.com URL:

http://zeus.xprt.com/˜ info/xprt.html Contact:

Jody Winston Type of support:

Phone, EMail and onsite. Special expertise:

Training, custom software, custom device drivers, custom application development. Sample prices:

Either by the hour starting at US$100 or a firm fixed price. Last modified:

February 19, 1997.

64 USA - Virginia

64.1 Internet Systems and Services

          6022 Rockton Court
          Centreville, VA 22020
          U.S.A.

Phone:

+1 703 222 4243 Fax:

+1 703 222 7320 EMail:

bass@silkroad.com URL:

http://www.silkroad.com Contact:

Tim Bass Type of support:

Phone, EMail and inhouse. Special expertise:

Large scale Internet systems, network security. Sample prices:

US$85/hour plus travel expenses (80 hours normal minimum). Last modified:

March 16, 1997.

65 USA - Washington

65.1 Cutting Edge Communications, Inc.

          P.O. Box 861
          Otis Orchards, WA 99027-0861
          U.S.A.

Phone:

+1 509 927 9541 Fax:

Used for emergency needs only. EMail:

roberth@cet.com URL:

http://www.cet.com Contact:

Robert Hanson Type of support:

Phone, EMail, remote network administration and onsite, Special expertise:

Programming, system installation and administration, Internet and Intranet connectivity, turn-key ISP solutions, LAN and WAN, security, systems design, engineering and integration, troubleshooting and training.

Sample prices:

US$75/hour to US$220/hour depending on client needs. Last modified:

March 21, 1997.

66 USA - Wisconsin

66.1 Couvares Consulting

     211 West Gilman St., Suite 3W
     Madison, WI 53703
     U.S.A.

Phone:

+1 608 256 6201 Fax:

+1 608 256 6201 EMail:

couvares@poboxes.com Contact:

Peter Couvares Type of support:

Phone, EMail, remote network administration and onsite. Special expertise:

Consulting, design, and implementation of customized systems and software solutions for free and commercial Unix environments.

Sample prices:

• US$60/hour for commercial clients

• US$20-60/hour sliding scale for personal and non-profit clients

Last modified:

February 21, 1997.